Theory and Applications of Categories, Vol. 14, No. 1, 2005, pp. 1–45.

CATEGORICAL STRUCTURES ENRICHED IN A QUANTALOID: CATEGORIES, DISTRIBUTORS AND FUNCTORS

ISAR STUBBE

Abstract. We thoroughly treat several familiar and less familiar definitions and results concerning categories, functors and distributors enriched in a base quantaloid Q. In analogy with V-category theory we discuss such things as adjoint functors, (pointwise) left Kan extensions, weighted (co)limits, presheaves and free (co)completion, Cauchy completion and Morita equivalence. With an appendix on the universality of the quantaloid Dist(Q) of Q-enriched categories and distributors.

1. Introduction

The theory of categories enriched in a symmetric monoidal closed category V is, by now, well known [Bénabou, 1963, 1965; Eilenberg and Kelly, 1966; Lawvere, 1973; Kelly, 1982]. For such a V with “enough” (co)limits the theory of V-categories, distributors and functors can be pushed as far as needed: it includes such things as (weighted) (co)limits in a V-category, V-presheaves on a V-category, Kan extensions of enriched functors, Morita theory for V-categories, and so on.

Monoidal categories are precisely one-object bicategories [Bénabou, 1967]. It is thus natural to ask how far V-category theory can be generalized to W-category theory, for W a general bicategory. But, whereas in V-category theory one usually assumes the symmetry of the tensor in V (which is essential for showing that V is itself a V-category with homobjects given by the right adjoint to tensoring), in working over a general bicategory W we will have to sacrifice this symmetry: tensoring objects in V corresponds to composing morphisms in W and in general it simply does not make sense for the composition g ∘ f of two arrows f, g to be “symmetric”.

On the other hand, we can successfully translate the notion of closedness of a monoidal category V to the more general setting of a bicategory W: ask that, for any object X of W and any arrow f : A [figure] B in W, both functors

− ∘ f : W(B, X) [figure] W(A, X) : x ↦ x ∘ f,  (1)
f ∘ − : W(X, A) [figure] W(X, B) : x ↦ f ∘ x  (2)

have respective right adjoints

{f, −} : W(A, X) [figure] W(B, X) : y ↦ {f, y},  (3)

Received by the editors 2004-05-13 and, in revised form, 2004-01-14.

Transmitted by Jiri Rosicky. Published on 2005-01-18.

2000 Mathematics Subject Classification: 06F07, 18B35, 18D20.

Key words and phrases: Quantales and quantaloids, enriched categories.

\begin{matrix}c \\ ◯ \end{matrix} Isar Stubbe, 2004. Permission to copy for private use granted.

[f, −] : W(X, B) [figure] W(X, A) : y ↦ [f, y].  (4)

Such a bicategory W is said to be closed. Some call an arrow such as {f, y} a (right) extension and [f, y] a (right) lifting (of y through f).

Finally, by saying that V has “enough limits and colimits” is in practice often meant that V has small limits and small colimits. In a bicategory W the analogue is straightforward: now ask for W to have in its hom-categories small limits and small colimits (i.e. W is locally complete and cocomplete).

So, to summarize, when trying to develop category theory over a base bicategory W rather than a base monoidal category V, it seems reasonable to work with a base bicategory which is closed, locally complete and locally cocomplete. Note that in such a bicategory W, due to its closedness, composition always distributes on both sides over colimits of morphisms:

f ∘ (colim_{i∈I} g_{i}) ≅ colim_{i∈I}(f ∘ g_{i}),  (5)
(colim_{j∈J} f_{j}) ∘ g ≅ colim_{j∈J}(f_{j} ∘ g).  (6)

That is to say, the local colimits are stable under composition. (But this does not hold in general for local limits!)

We will focus on a special case of these closed, locally complete and locally cocomplete bicategories: namely, we study such bicategories whose hom-categories are moreover small and skeletal. Thus the hom-categories are simply complete lattices. We will write the local structure as an order, and local limits and colimits of morphisms as their infimum, resp. supremum—so for arrows with same domain and codomain we have things like f ≤ f′, ⋁_{i∈I} f_{i}, ⋀_{j∈J} g_{j}, etc. In particular (5) and (6) become

f ∘ (⋁_{i} g_{i}) = ⋁_{i}(f ∘ g_{i}),  (7)
(⋁_{j} f_{j}) ∘ g = ⋁_{j}(f_{j} ∘ g).  (8)

The adjoint functor theorem says that the existence of the adjoints (3) and (4) to the composition functors (1) and (2) (not only implies but also) is implied by their distributing over suprema of morphisms as in (7) and (8). Such bicategories – whose hom-categories are complete lattices and whose composition distributes on both sides over arbitrary suprema – are called quantaloids. A one-object quantaloid is a quantale^{1}. So a quantaloid Q is a Sup-enriched category (and a quantale is monoid in Sup).

We will argue below that “V-category theory” can be generalized to “Q-category theory”, where now Q is a quantaloid. This is a particular case of the theory of “W-category theory” as pioneered by [Bénabou, 1967; Walters, 1981; Street, 1983a]. But we feel that this particular case is also of particular interest: many examples of bicategory-enriched categories are really quantaloid-enriched. Also, without becoming trivial, quantaloidenrichment often behaves remarkably better than general bicategory-enrichment: essentially because all diagrams of 2-cells in a quantaloid commute. So, for example, when calculating extensions and liftings in a quantaloid, we need not keep track of the universal 2-cells; or when defining a Q-category, we need not specify its composition- and identity-2-cells nor impose coherence axioms; also the composition of distributors simplifies a great deal, for we need not compute complicated colimits of 1-cells in a general bicategory but simply suprema of arrows in a quantaloid; and so on. In our further study of categorical structures enriched in a quantaloid we rely heavily on the basic theory of Q-categories, distributors and functors; however, often we could not find an appropriate reference for one or another basic fact. With this text we wish to provide such a reference.

1It was C. Mulvey [1986] who introduced the word ‘quantale’ in his work on (non-commutative)

C^{∗}-algebras as a contraction of ‘quantum’ and ‘locale’.

A first further study, in [Stubbe, 2004a], is concerned with “variation and enrichment”, in the sense of [Betti et al., 1983; Gordon and Power, 1997, 1999]. This is a development of the notion of weighted colimit in a Q-enriched category; in particular, a tensored and cotensored Q-category has all weighted colimits (as in section 5 below) if and only if it has all conical colimits, if and only if its underlying order has suprema of objects of the same type. This allows for a detailed analysis of the biequivalence between Q-modules and cocomplete Q-categories.

The subject of [Stubbe, 2004b] is that of presheaves on Q-semicategories (“categories without units”), along the lines of [Moens et al., 2002]; it generalizes the theory of regular modules on an R-algebra without unit, for R a commutative ring (with unit). The point is that certain “good properties” of the Yoneda embedding for Q-categories (see section 6 below) are no longer valid for Q-semicategories—for example, a presheaf on a Q-semicategory is not canonically the weighted colimit of representables. Enforcing precisely this latter condition defines what is called a “regular presheaf” on a Q-semicategory; there is an interesting theory of “regular Q-semicategories”.

[Borceux and Cruciani, 1998] gives an elementary description of ordered objects in the topos of sheaves on a locale Ω; this turns out to be all about enriched Ω-semicategories that admit an appropriate Cauchy completion. In [Stubbe, 2004c] we describe more generally the theory of Q-semicategories that admit a well-behaved Cauchy completion (based on the material in section 7): these Cauchy complete semicategories are what we believe to be the correct notion of “ordered sheaves on a quantaloid Q”, and we simply call them ‘Q-orders’. Such Q-orders can equivalently be described as categories enriched in the split-idempotent completion of Q, and so provide a missing link between [Walters, 1981] and [Borceux and Cruciani, 1998]. We hope that the theory of Q-orders, as a theory of “(ordered) sheaves on a non-commutative topology”, will provide new grounds for defining notions of spectra for non-commutative rings and C^{∗}-algebras—an issue which was at the origin of the theory of quantales and quantaloids!

Acknowledgment. I thank Francis Borceux and Ross Street for their generosity, not only in mathematics, but simply in general.

2. Quantaloids

A sup-lattice is an antisymmetrically ordered set^{2} (X, ≤) for which every subset X′ ⊆ X has a supremum ⋁ X′ in X. A morphism of sup-lattices f: (X, ≤) [figure] (Y, ≤) is a map f: X [figure] Y that preserves arbitrary suprema. It is well-known that sup-lattices and supmorphisms constitute a symmetric monoidal closed category Sup.

2.1. Definition. A quantaloid Q is a Sup-enriched category. A homomorphism F: Q [figure] Q′ of quantaloids is a Sup-enriched functor.

In principle we do not mind a quantaloid having a proper class of objects. Thus quantaloids and homomorphisms form an illegitimate category QUANT; small quantaloids define a (true) subcategory Quant. A quantaloid with one object is often thought of as a monoid in Sup, and is called a quantale.

In elementary terms, a quantaloid Q is a category whose hom-sets are actually suplattices, in which composition distributes on both sides over arbitrary suprema of morphisms. In the same vein, a homomorphism F: Q [figure] Q′ is a functor of (the underlying) categories that preserves arbitrary suprema of morphisms.

For arrows f: A [figure] B, g: B [figure] C, (h_{i}: X [figure] Y)_{i∈I} in a quantaloid Q we use notations like g ∘ f: A [figure] C for composition, and ⋁_{i} h_{i}: X [figure] Y and h_{i} ≤ h_{j}: X [figure] Y for its local structure; Q(A, B) is the hom-lattice of arrows from A to B. The identity arrow on an object A ∈ Q is written 1_{A}: A [figure] A. The bottom element of a sup-lattice Q(A, B) will typically be denoted by 0_{A,B}. With these notations for identity and bottom, we can write a “Kronecker delta”

δ_{A,B}: A [figure] B = \{ \begin{matrix}1_{A}: A [figure] A & if A = B, \\ 0_{A,B}: A [figure] B & otherwise. \end{matrix}

As we may consider Sup to be a “simplified version of Cat”, we may regard quantaloids as “simplified bicategories”. Notably, a quantaloid Q has small hom-categories with stable local colimits, and therefore it is closed. Considering morphisms f: A [figure] B and g: B [figure] C we note the respective adjoints to composition in Q as − ∘ f ⊣ {f, −} and g ∘ − ⊣ [g, −]; that is to say, for any h: A [figure] C we have that g ∘ f ≤ h iff f ≤ [g, h] iff g ≤ {f, h}. Note furthermore that every diagram of 2-cells in a quantaloid trivially commutes.

There are a couple of lemmas involving [−, −] and {−, −}, that hold in any quantaloid, upon which we rely quite often. Let us give a short overview.

2.2. Lemma. We work in a quantaloid Q.

1. For f: A [figure] B, g: B [figure] C and h: A [figure] C, [g, h] = ⋁{x: A [figure] B | g ∘ x ≤ h} and {f, h} = ⋁{y: B [figure] C | y ∘ f ≤ h}.

2. The following are equivalent:

2By an “ordered set” we mean a set endowed with a transitive, reflexive relation; it is what is often called a “preordered set”. We will be explicit when we mean one such relation that is moreover antisymmetric, i.e. a “partial order”.

(a) f ⊣ g: B [figure] A in Q (i.e. 1_{A} ≤ g ∘ f and f ∘ g ≤ 1_{B});

(b) (f ∘ −) ⊣ (g ∘ −): Q(−, B) [figure] Q(−, A) in Sup;

(c) (− ∘ g) ⊣ (− ∘ f): Q(B, −) [figure] Q(A, −) in Sup;

(d) (g ∘ −) = [f, −]: Q(−, B) [figure] Q(−, A) in Sup;

(e) (− ∘ f) = {g, −}: Q(B, −) [figure] Q(A, −) in Sup.

3. An arrow f: A [figure] B has a right adjoint if and only if [f, 1_{B}] ∘ f = [f, f]; in this case the right adjoint to f is [f, 1_{B}]. Dually, g: B [figure] A in Q has a left adjoint if and only if g ∘ {g, 1_{B}} = {g, g}; in this case the left adjoint to g is {g, 1_{B}}.

4. For f, f′: A [figure] B and g, g′: B [figure] A such that f ⊣ g and f′ ⊣ g′, f ≤ f′ if and only if g′ ≤ g.

5. Any f: A [figure] B induces, for every X ∈ Q_{0}, an adjunction

[−, f] ⊣ {−, f}: Q(X, B) [figure] Q(A, X)^{op}.

By Q(A, X)^{op} is meant the sup-lattice Q(A, X) with opposite order.

6. For arrows as in the diagrams below, we have the identity [f, [g, h]] = [g ∘ f, h], its dual {k, {l, m}} = {k ∘ l, m} and the self-dual [x, {y, z}] = {y, [x, z]}.

[figure: B; g; C; X; k; Y; l; m; I; z; y; J; A; f; D; h; W; Z; H; K; x]

7. For arrows as in the diagrams below, we have the identity [h, g] ∘ [g, f] ≤ [h, f] and its dual {l, m} ∘ {k, l} ≤ {k, m}. Also 1_{dom(f)} ≤ [f, f], and dually 1_{cod(f)} ≤ {f, f}, hold.

[figure: A; f; T; S; k; l; m; X; B; g; C; h; Z; Y]

Sup may also be thought of as an “infinitary version of Ab”; a quantaloid is then like a “ring(oid) with infinitary (and idempotent) sum”. This point of view helps explain some of the terminology below—especially when we talk about “matrices with elements in a quantaloid Q” in the appendix.

2.3. Example. Any locale Ω is a (very particular) quantale. Further on we shall denote by 2 the two-element boolean algebra. Given a locale Ω, [Walters, 1981] uses the quantaloid of relations in Ω: the objects of Rel(Ω) are the elements of Ω, the hom-lattices are given by Rel(Ω)(u, v) = {w ∈ Ω | w ≤ u∧v} with order inherited from Ω, composition is given by infimum in Ω, and the identity on an object u is u itself. This quantaloid is very particular: it equals its opposite.

2.4. Example. The extended non-negative reals [0, +∞] form a (commutative) quantale for the opposite order and addition as binary operation [Lawvere, 1973].

2.5. Example. The ideals in a commutative ring R form a (commutative) quantale. When R is not commutative, the two-sided ideals still form a (commutative) quantale. But, as G. Van den Bossche [1995] points out (but she credits B. Lawvere for this idea), there is also quite naturally a “quantaloid of ideals” containing a lot more information than just the two-sided ideals. Denoting Q_{R} for this structure, define that it has two objects, 0 and 1; the hom-sup-lattices are

Q_{R}(0, 0) = additive subgroups of R which are Z(R)-modules, Q_{R}(0, 1) = left-sided ideals of R, Q_{R}(1, 0) = right-sided ideals of R, Q_{R}(1, 1) = two-sided ideals of R,

with sum of additive subgroups as supremum:

∑_{k∈K} I_{k} = {finite sums of elements in ⋃_{k∈K} I_{k}}.

Composition in Q_{R} is the multiplication of additive subgroups, as in

I ∘ J = {finite sums i_{1}j_{1} + ... + i_{n}j_{n} with all i_{k} ∈ I and all j_{k} ∈ J};

and the identity arrow on 1 is R and that on 0 is Z(R). (By Z(R) we denote the center of R). For a commutative R, this quantaloid Q_{R} is equivalent as Sup-category to the quantale of ideals in R.

More examples can be found in the literature, e.g. [Rosenthal, 1996].

3. Three basic definitions

In the following Q always stands for a quantaloid, and Q_{0} for its (possibly large) class of objects. A Q-typed set X is a (small) set X together with a mapping t: X [figure] Q_{0}: x ↦ tx sending each element x ∈ X to its type tx ∈ Q_{0}. The notation with a “t” for the types of elements in a Q-typed set is generic; i.e. even for two different Q-typed sets X and Y, the type of x ∈ X is written tx, and that of y ∈ Y is ty. A Q-typed set X is just a way of writing a set-indexed family (tx)_{x∈X} of objects (i.e. a small discrete diagram) in Q. If Q is a small quantaloid, then a Q-typed set is an object of the slice category Set/Q_{0}.

3.1. Definition. A Q-enriched category (or Q-category for short) 𝔸 consists of

- objects: a Q-typed set 𝔸_{0},

- hom-arrows: for all a, a′ ∈ 𝔸_{0}, an arrow 𝔸(a′, a): ta [figure] ta′ in Q,

satisfying

- composition-inequalities: for all a, a′, a′′ ∈ 𝔸_{0}, 𝔸(a′′, a′) ∘ 𝔸(a′, a) ≤ 𝔸(a′′, a) in Q,

- identity-inequalities: for all a ∈ 𝔸_{0}, 1_{ta} ≤ 𝔸(a, a) in Q.

3.2. Definition. A distributor Φ: 𝔸 [figure] 𝔹 between two Q-categories is given by

- distributor-arrows: for all a ∈ 𝔸_{0}, b ∈ 𝔹_{0}, an arrow Φ(b, a): ta [figure] tb in Q

satisfying

- action-inequalities: for all a, a′ ∈ 𝔸_{0}, b, b′ ∈ 𝔹_{0}, 𝔹(b′, b) ∘ Φ(b, a) ≤ Φ(b′, a) and Φ(b, a) ∘ 𝔸(a, a′) ≤ Φ(b, a′) in Q.

3.3. Definition. A functor F: 𝔸 [figure] 𝔹 between Q-categories is

- object-mapping: a map F: 𝔸_{0} [figure] 𝔹_{0}: a ↦ F a

satisfying

- type-equalities: for all a ∈ 𝔸_{0}, ta = t(F a) in Q,

- action-inequalities: for all a, a′ ∈ 𝔸_{0}, 𝔸(a′, a) ≤ 𝔹(F a′, F a) in Q.

None of these definitions requires any of the usual diagrammatic axioms (associativity and identity axioms for the composition in a category, coherence of the action on a distributor with the composition in its (co)domain category, the functoriality of a functor) simply because those conditions, which require the commutativity of certain diagrams of 2-cells in the base quantaloid Q, hold trivially! It is therefore a property of, rather than an extra structure on, a given set 𝔸_{0} of elements with types in Q together with homarrows 𝔸(a′, a): ta′ [figure] ta (for a, a′ ∈ 𝔸_{0}) to be a Q-category; in other words, if these data determine a Q-category, then they do so in only one way. Similar for distributors and functors: it is a property of a given collection of arrows Φ(b, a): ta [figure] tb whether or not it determines a distributor between Q-categories 𝔸 and 𝔹, as it is a property of an object mapping F: 𝔸_{0} [figure] 𝔹_{0} whether or not it determines a functor.

Our Q-categories are by definition small: they have a set of objects. As a consequence, the collection of distributors between two Q-categories is always a small set too, and so is the collection of functors between two Q-categories. (However, we will soon run into size-related trouble: our base quantaloid Q having a proper class of objects will conflict with the Q-categories being small, in particular in matters related to cocompleteness of Q-categories. When doing categorical algebra over Q we will therefore suppose that Q is small too. But this hypothesis is not necessary at this point.)

Note that for a distributor Φ: 𝔸 [figure] 𝔹, for all a ∈ 𝔸 and b ∈ 𝔹,

Φ(b, a) = 1_{tb} ∘ Φ(b, a) ≤ 𝔹(b, b) ∘ Φ(b, a) ≤ ⋁_{b′∈𝔹}𝔹(b, b′) ∘ Φ(b′, a) ≤ Φ(b, a),

that is to say, ⋁_{b′∈𝔹} 𝔹(b, b′) ∘ Φ(b′, a) = Φ(b, a); ⋁_{a′∈𝔸} Φ(b, a′) ∘ 𝔸(a′, a) = Φ(b, a) is analogous. A Q-category 𝔸 is itself a distributor from 𝔸 to 𝔸; and the identities above become, for all a, a′′ ∈ 𝔸, ⋁_{a′∈𝔸} 𝔸(a′′, a′) ∘ 𝔸(a′, a) = 𝔸(a′′, a). These identities allow to say, in 3.4, with a suitable definition for the composition of distributors, that “𝔸 is the identity distributor on 𝔸”.

There is a notational issue that we should comment on. We have chosen to write the composition of arrows in a base quantaloid Q “from right to left”: the composite of f: X [figure] Y and g: Y [figure] Z is g ∘ f: X [figure] Z. Therefore we have chosen to write the homarrows in a Q-enriched category 𝔸 also “from right to left”: for two objects a, a′ ∈ 𝔸, the hom-arrow 𝔸(a′, a) goes from ta to ta′. Doing so it is clear that, for example, the composition-inequality in 𝔸 is written 𝔸(a′′, a′) ∘ 𝔸(a′, a) ≤ 𝔸(a′′, a), with the pivot a′ nicely in the middle, which we find very natural. Our notational conventions are thus basically those of R. Street’s seminal paper [1983a]; other authors have chosen other notations.

The next proposition displays the calculus of Q-categories and distributors.

3.4. Proposition. Q-categories are the objects, and distributors the arrows, of a quantaloid Dist(Q) in which

- the composition Ψ ⊗_{𝔹} Φ: 𝔸 [figure] ℂ of two distributors Φ: 𝔸 [figure] 𝔹 and Ψ: 𝔹 [figure] ℂ has as

distributor-arrows, for a ∈ 𝔸_{0} and c ∈ ℂ_{0},

(Ψ ⊗_{𝔹} Φ)(c, a) = ⋁_{b∈𝔹_{0}} Ψ(c, b) ∘ Φ(b, a);

- the identity distributor on a Q-category 𝔸 has as distributor-arrows precisely the hom-arrows of the category 𝔸 itself, so we simply write it as 𝔸: 𝔸 [figure] 𝔸;

- the supremum ⋁_{i∈I} Φ_{i}: 𝔸 [figure] 𝔹 of given distributors (Φ_{i}: 𝔸 [figure] 𝔹)_{i∈I} is calculated elementwise, thus its distributor-arrows are, for a ∈ 𝔸_{0} and b ∈ 𝔹_{0},

(⋁_{i∈I} Φ_{i})(b, a) = ⋁_{i∈I} Φ_{i}(b, a).

The proof of the fact that the data above define a quantaloid is straightforward. Actually, Dist(Q) is a universal construction on Q in QUANT: there is a fully faithful homomorphism of quantaloids

Q [figure] Dist(Q): (f: A [figure] B) ↦ ((f): ∗_{A} [figure]∗_{B} )  (9)

sending an object A ∈ Q to the Q-category with only one object, say ∗, of type t∗ = A, and hom-arrow 1_{A}, and a Q-arrow f: A [figure] B to the distributor (f): ∗_{A} [figure] ∗_{B} whose single element is f. This turns out to be the universal direct-sum-and-split-monad completion of Q in QUANT. The appendix gives details. (Even for a small base quantaloid Q, Dist(Q) has a proper class of objects; so large quantaloids arise as universal constructions on small ones. Therefore we do not wish to exclude large quantaloids a priori.)

Since Dist(Q) is a quantaloid, it is in particular closed; the importance of this fact cannot be overestimated. Let for example Θ: 𝔸 [figure] ℂ and Ψ: 𝔹 [figure] ℂ be distributors between Q-categories, then [Ψ, Θ]: 𝔸 [figure] 𝔹 is the distributor with distributor-arrows, for a ∈ 𝔸_{0} and b ∈ ℂ_{0},

[Ψ, Θ](b, a) = ⋀_{c∈ℂ_{0}} [Ψ(c, b), Θ(c, a)],  (10)

where the liftings on the right are calculated in Q. A similar formula holds for {−, −}. The category of Q-categories and functors is the obvious one.

3.5. Proposition. Q-categories are the objects, and functors the arrows, of a category Cat(Q) in which

- the composition G∘F: 𝔸 [figure] ℂ of two functors F: 𝔸 [figure] 𝔹 and G: 𝔹 [figure] ℂ is determined by the composition of object maps G ∘ F: 𝔸_{0} [figure] ℂ_{0}: a ↦ G(F (a));

- the identity functor 1_{𝔸}: 𝔸 [figure] 𝔸 on a Q-category 𝔸 is determined by the identity object map 1_{𝔸}: 𝔸_{0} [figure] 𝔸_{0}: a ↦ a.

Every functor between Q-categories induces an adjoint pair of distributors, and the resulting inclusion of the functor category in the distributor category – although straightforward – is a key element for the development of the theory of Q-enriched categories.

3.6. Proposition. For Q-categories and functors F: 𝔸 [figure] 𝔹 and G: ℂ [figure] 𝔹, the Q-arrows 𝔹(Gc, F a): ta [figure] tc, one for each (a, c) ∈ 𝔸_{0} × ℂ_{0}, determine a distributor^{3} 𝔹(G−, F −): 𝔸 [figure] ℂ. For any functor F: 𝔸 [figure] 𝔹 the distributors 𝔹(1_{𝔹}−, F −): 𝔸 [figure] 𝔹 and 𝔹(F −, 1_{𝔹}−): 𝔹 [figure] 𝔸 are adjoint in the quantaloid Dist(Q): 𝔹(1_{𝔹}−, F −) ⊣ 𝔹(F −, 1_{𝔹}−).

Proof. For any a, a′ ∈ 𝔸_{0} and c, c′ ∈ ℂ_{0}, ℂ(c′, c) ∘ 𝔹(Gc, F a) ∘ 𝔸(a, a′) ≤ 𝔹(Gc′, Gc) ∘ 𝔹(Gc, F a) ∘ 𝔹(F a, F a′) ≤ 𝔹(Gc′, F a) by functoriality of F and G and composition in 𝔹. So 𝔹(G−, F −) is a distributor from 𝔸 to 𝔹.

To see that 𝔹(1_{𝔹}−, F −) ⊣ 𝔹(F −, 1_{𝔹}−) in Dist(Q), we must check two inequalities: the unit of the adjunction is due to the composition in 𝔹 and functoriality of F, 𝔹(F −, 1_{𝔹}−)⊗_{𝔹} 𝔹(1_{𝔹}−, F −) = 𝔹(F −, F −) ≥ 𝔸(−, −); the counit follows from the fact that {F a | a ∈ 𝔸_{0}} ⊆ 𝔹_{0} and – again – composition in 𝔹, 𝔹(1_{𝔹}−, F −) ⊗_{𝔸} 𝔹(F −, 1_{𝔹}−) ≤ 𝔹(1_{𝔹}−, −) ⊗_{𝔹} 𝔹(−, 1_{𝔹}−) = 𝔹(−, −).

3There is a converse: for any distributor Φ: 𝔸 [figure] ℂ there is a – universal – way in which Φ = 𝔹(G−, F −) for certain functors F: 𝔸 [figure] 𝔹 and G: ℂ [figure] 𝔹.

In the following we use the abbreviated notations 𝔹(−, F −) = 𝔹(1_{𝔹}−, F −) and 𝔹(F −, −) = 𝔹(F −, 1_{𝔹}−).

3.7. Proposition. Sending a functor to the left adjoint distributor that it induces, as in

Cat(Q) [figure] Dist(Q): (F: 𝔸 [figure] 𝔹) ↦ (𝔹(−, F −): 𝔸 [figure] 𝔹),

is functorial. Sending a functor to the right adjoint determines a similar contravariant functor.

Proof. Trivially an identity functor 1_{𝔸}: 𝔸 [figure] 𝔸 is mapped onto the identity distributor 𝔸(−, 1_{𝔸}−) = 𝔸(−, −): 𝔸 [figure] 𝔸. And if F: 𝔸 [figure] 𝔹 and G: 𝔹 [figure] ℂ, then we want ℂ(−, G ∘ F −) = ℂ(−, G−) ⊗_{𝔹} 𝔹(−, F −) but also this holds trivially: read the right hand side of the equation as the action of 𝔹 on ℂ(−, G−).

The category Cat(Q) inherits the local structure from the quantaloid Dist(Q) via the functor Cat(Q) [figure] Dist(Q): we put, for two functors F, G: 𝔸 [figure] 𝔹,

F ≤ G ⟺ 𝔹(−, F −) ≤ 𝔹(−, G−) ( ⟺ 𝔹(G−, −) ≤ 𝔹(F −, −)).

Thus every hom-set Cat(Q)(𝔸, 𝔹) is (neither antisymmetrically nor cocompletely) ordered, and composition in Cat(Q) distributes on both sides over the local order: Cat(Q) is a 2category, and the functor in 3.7 is a 2-functor which is the identity on objects, “essentially faithful” (but not full), and locally fully faithful. (The contravariant version of this functor reverses the local order!)

In general the “opposite” of a Q-category 𝔸 is not again a Q-category, but rather a Q^{op}-category: of course 𝔸^{op} is defined to have the same Q-typed set of objects as 𝔸, but the hom-arrows are reversed: for objects a, a′ put 𝔸^{op}(a′, a) = 𝔸(a, a′). Similarly, for a distributor Φ: 𝔸 [figure] 𝔹 between Q-categories we may define an opposite distributor between the opposite categories over the opposite base—but this distributor will go in the opposite direction: Φ^{op}: 𝔹^{op} [figure] 𝔸^{op} is defined by Φ^{op}(a, b) = Φ(b, a). For Φ ≤ Ψ: 𝔸 [figure] 𝔹 in Dist(Q) it is quite obvious that Φ^{op} ≤ Ψ^{op}: 𝔹^{op} [figure] 𝔸^{op} in Dist(Q^{op}). Finally, for a functor F: 𝔸 [figure] 𝔹 between Q-categories, the same object mapping a ↦ F a determines an arrow F^{op}: 𝔸^{op} [figure] 𝔹^{op} of Cat(Q^{op}). But if F ≤ G: 𝔸 [figure] 𝔹 in Cat(Q) then G^{op} ≤ F^{op}: 𝔸^{op} [figure] 𝔹^{op} in Cat(Q^{op}). It is obvious that applying the “op” twice, always gives back the original structure.

3.8. Proposition. “Taking opposites” determines isomorphisms

Dist(Q) ≅ Dist(Q^{op})^{op} and Cat(Q) ≅ Cat(Q^{op})^{co}

of 2-categories (where the “co” means: reversing order in the homs).

These isomorphisms allow us to “dualize” all notions and results concerning Q-enriched categories. For example, the dual of ‘left Kan extension’ in Cat(Q) is ‘right Kan extension’, and whereas the former help to characterize left adjoints in Cat(Q), the latter do the same for right adjoints. As another example, the dual of ‘weighted colimit’ in a Q-category, is ‘weighted limit’; the former are ‘preserved’ by all left adjoint functors, the latter by right adjoint ones. Some notions are self-dual, like ‘equivalence’ of categories, or ‘Cauchy complete’ category.

3.9. Example. In a quantaloid Q, the arrows whose codomain is some object Y are the objects of a Q-category that we will denote – anticipating 6.1 – by PY: put t(f: X [figure] Y) = X and PY (f′, f) = [f′, f]. Similarly – and again anticipating further results – we denote by P^{†}X the Q-category whose objects are Q-arrows with domain X, with types t(f: X [figure] Y) = Y, and P^{†}X(f′, f) = {f, f′}.

3.10. Example. Recall that 2 is the 2-element Boolean algebra; 2-categories are orders, distributors are ideal relations, and functors are order-preserving maps.

3.11. Example. Consider a sheaf F on a locale Ω; it determines a Rel(Ω)-category 𝔽 whose objects are the partial sections of F, the type of a section s being the largest u ∈ Ω on which it is defined, and whose hom-arrows 𝔽(s′, s) are, for sections s, s′ of types u, u′, the largest v ≤ u ∧ u′ on which (restrictions of) s and s′ agree [Walters, 1981].

3.12. Example. A category enriched in the quantale [0, +∞] (cf. 2.4) is a “generalized metric space” [Lawvere, 1973]: the enrichment itself is a binary distance function taking values in the positive reals. In particular is the composition-inequality in such an enriched category the triangular inequality. A functor between such generalized metric spaces is a distance decreasing application. (These metric spaces are “generalized” in that the distance function is not symmetric, that the distance between two points being zero does not imply their being identical, and that the distance between two points may be infinite.)

3.13. Example. A (not necessarily commutative) ring R determines a Q_{R}-enriched category (with Q_{R} as in 2.5): denoting it as Comm_{R}, its objects of type 0 are the elements of R, its objects of type 1 are the elements of Z(R), and hom-arrows are given by commutators: Comm(r, s) = {x ∈ R | rx = xs}.

4. Some direct consequences

Underlying orders

For an object A of a quantaloid Q, denote by ∗_{A} the one-object Q-category whose homarrow is the identity 1_{A}. Given a Q-category 𝔸, the set {a ∈ 𝔸_{0} | ta = A} is in bijection with Cat(Q)(∗_{A}, 𝔸): any such object a determines a “constant” functor Δa: ∗_{A} [figure] 𝔸; and any such functor F: ∗_{A} [figure] 𝔸 “picks out” an object a ∈ 𝔸. We may thus order the objects of 𝔸 by ordering the corresponding constant functors; we will speak of the underlying order (𝔸_{0}, ≤) of the Q-category 𝔸. Explicitly, for two objects a, a′ ∈ 𝔸_{0} we have that a′ ≤ a if and only if A := ta = ta′ and for all x ∈ 𝔸_{0}, 𝔸(x, a′) ≤ 𝔸(x, a) in Q, or equivalently 𝔸(a′, x) ≥ 𝔸(a, x), or equivalently 1_{A} ≤ 𝔸(a′, a). Whenever two objects of 𝔸 are equivalent in 𝔸’s underlying order (a ≤ a′ and a′ ≤ a) then we say that they are isomorphic objects (and write a ≅ a′).

4.1. Example. In the Q-category PY, whose objects are Q-arrows with codomain Y and whose hom-arrows are given by [−, −], the underlying order coincides with the local order in Q. More precisely, for f, g: X [figure] Y it is the same to say that f ≤ g in (PY)_{0} as to say that f ≤ g in Q(X, Y).

It is immediate that F ≤ G: 𝔸 [figure] 𝔹 in Cat(Q) if and only if, for all a ∈ 𝔸_{0}, F a ≤ Ga in the underlying order of 𝔹. This says that the local structure in the 2-category Cat(Q) is “pointwise order”. Equivalently we could have written that F ≤ G if and only if 1_{ta} ≤ 𝔹(F a, Ga) for all a ∈ 𝔸_{0}, which exhibits the resemblance with the usual notion of “enriched natural transformation”.

Adjoints and equivalences

An arrow F: 𝔸 [figure] 𝔹 is left adjoint to an arrow G: 𝔹 [figure] 𝔸 in Cat(Q) (and G is then right adjoint to F), written F ⊣ G, if 1_{𝔸} ≤ G ∘ F and F ∘ G ≤ 1_{𝔹}. Due to the 2-category Cat(Q) being locally ordered, we need not ask any of the usual triangular coherence diagrams. The unicity of adjoints in the quantaloid Dist(Q) and the locally fully faithful Cat(Q) [figure] Dist(Q) allow for the following equivalent expression.

4.2. Proposition. F: 𝔸 [figure] 𝔹 is left adjoint to G: 𝔹 [figure] 𝔸 in Cat(Q) if and only if 𝔹(F −, −) = 𝔸(−, G−): 𝔹 [figure] 𝔸 in Dist(Q).

Further, F: 𝔸 [figure] 𝔹 is an equivalence in Cat(Q) if there exists a G: 𝔹 [figure] 𝔸 such that G ∘ F ≅ 1_{𝔸} and F ∘ G ≅ 1_{𝔹} (in which case also G is an equivalence). Again because Cat(Q) is locally ordered, this is the same as saying that F is both left and right adjoint to some G. Again the functor Cat(Q) [figure] Dist(Q) gives equivalent expressions.

4.3. Proposition. Given functors F: 𝔸 [figure] 𝔹 and G: 𝔹 [figure] 𝔸 constitute an equivalence in Cat(Q) if and only if 𝔹(−, F −): 𝔸 [figure] 𝔹 and 𝔸(−, G−): 𝔹 [figure] 𝔸 constitute an isomorphism in Dist(Q), if and only if 𝔹(F −, −): 𝔹 [figure] 𝔸 and 𝔸(G−, −): 𝔸 [figure] 𝔹 constitute an isomorphism in Dist(Q).

For ordinary categories and functors it is well known that the equivalences are precisely the fully faithful functors which are essentially surjective on objects. This holds for Qenriched categories too: say that a functor F: 𝔸 [figure] 𝔹 is fully faithful if

∀ a, a′ ∈ 𝔸_{0} : 𝔸(a′, a) = 𝔹(F a′, F a) in Q,

and that is essentially surjective on objects whenever

∀ b ∈ 𝔹_{0}, ∃ a ∈ 𝔸_{0} : F a ≅ b in 𝔹.

In fact, F is fully faithful if and only if the unit of the adjunction 𝔹(−, F −) ⊣ 𝔹(F −, −) in Dist(Q) saturates to an equality; and if F is essentially surjective on objects then necessarily the co-unit of this adjunction saturates to an equality (but a functor F: 𝔸 [figure] 𝔹 for which the co-unit of the induced adjunction is an equality – which is sometimes said to have a “dense image” – is not necessarily essentially surjective on objects).

4.4. Proposition. An arrow F: 𝔸 [figure] 𝔹 in Cat(Q) is an equivalence if and only if it is fully faithful and essentially surjective on objects.

Proof. Suppose that F is an equivalence, with inverse equivalence G. Then, by functoriality of F and G, and since G∘F ≅ 1_{𝔸}, 𝔸(a′, a) ≤ 𝔹(F a′, F a) ≤ 𝔸(GF a′, GF a) = 𝔸(a′, a) for all a, a′ ∈ 𝔸_{0}. Further, using F ∘G ≅ 1_{𝔹}, it follows that b ≅ F Gb, and Gb ∈ 𝔸_{0} as required. Conversely, supposing that F is fully faithful and essentially surjective on objects, choose – using the essential surjectivity of F – for any b ∈ 𝔹_{0} one particular object Gb ∈ 𝔸_{0} for which F Gb ≅ b. Then the object mapping 𝔹_{0} [figure] 𝔸_{0}: b ↦ Gb is fully faithful itself (and therefore also functorial) due to the fully faithfulness of F: 𝔹(b′, b) = 𝔹(F Gb′, F Gb) = 𝔸(Gb′, Gb) for all b, b′ ∈ 𝔹_{0}. It is the required inverse equivalence.

The following is well known for ordinary categories, and holds for Q-categories too; it will be useful further on.

4.5. Proposition. Suppose that F ⊣ G: 𝔹 [figure] 𝔸 in Cat(Q). Then F is fully faithful if and only if G ∘ F ≅ 1_{𝔸}, and G is fully faithful if and only if F ∘ G ≅ 1_{𝔹}. If moreover G ⊣ H: 𝔸 [figure] 𝔹 in Cat(Q) then F is fully faithful if and only if H is fully faithful.

Proof. F ⊣ G in Cat(Q) implies 𝔹(F −, F −) = 𝔸(−, G ∘ F −) in Dist(Q); and F is fully faithful if and only if 𝔹(F −, F −) = 𝔸(−, −) in Dist(Q). So, obviously, F is fully faithful if and only if 𝔸(−, G ∘ F −) = 𝔸(−, −) in Dist(Q), or equivalently, G ∘ F ≅ 1_{𝔸} in Cat(Q). Likewise for the fully faithfulness of G. Suppose now that F ⊣ G ⊣ H, then 𝔸(G∘F −, −) = 𝔹(F −, H−) = 𝔹(−, G∘H−), which implies that 𝔸(G∘F −, −) = 𝔸(−, −) if and only if 𝔹(−, G ∘ H−) = 𝔸(−, −) so the result follows.

Left Kan extensions

Given Q-categories and functors F: 𝔸 [figure] 𝔹 and G: 𝔸 [figure] ℂ, the left Kan extension of F along G is – in so far it exists – a functor K: ℂ [figure] 𝔹 such that K ∘ G ≥ F in a universal way: whenever K′: ℂ [figure] 𝔹 satisfies K′ ∘ G ≥ F, then K′ ≥ K. If the left Kan extension of F along G exists, then it is essentially unique; we denote it by ⟨F, G⟩. So ⟨F, G⟩: ℂ [figure] 𝔹 is the reflection of F ∈ Cat(Q)(𝔸, 𝔹) along

− ∘ G: Cat(Q)(ℂ, 𝔹) [figure] Cat(Q)(𝔸, 𝔹): H ↦ H ∘ G.

Left adjoint functors may be characterized in terms of left Kan extensions; this uses the idea of functors that preserve Kan extensions. Suppose that the left Kan extension ⟨F, G⟩: ℂ [figure] 𝔹 exists; then a functor F′: 𝔹 [figure] 𝔹′ is said to preserve ⟨F, G⟩ if ⟨F′ ∘ F, G⟩ exists and is isomorphic to F′ ∘ ⟨F, G⟩. And ⟨F, G⟩ is absolute if it is preserved by any F′: 𝔹 [figure] 𝔹′.

4.6. Proposition. For a functor F: 𝔸 [figure] 𝔹 between Q-categories the following are equivalent:

1. F has a right adjoint; 2. ⟨1_{𝔸}, F ⟩ exists and is absolute;

3. ⟨1_{𝔸}, F ⟩ exists and is preserved by F itself.

In this case, ⟨1_{𝔸}, F ⟩ is the right adjoint of F.

Proof. For (1⇒2), suppose that F ⊣ G; we shall prove that G is the absolute left Kan extension of 1_{𝔸} along F. The unit of the adjunction already says that G ∘ F ≥ 1_{𝔸}; and if K: 𝔹 [figure] 𝔸 is another functor such that K ∘ F ≥ 1_{𝔸}, then necessarily K = K ∘ 1_{𝔹} ≥ K ∘ (F ∘ G) = (K ∘ F) ∘ G ≥ 1_{𝔸} ∘ G = G, using now the unit of the adjunction. So indeed G ≅ ⟨1_{𝔸}, F ⟩. Let now F′: 𝔸 [figure] 𝔹′ be any functor; we claim that F′ ∘ G is the left Kan extension of F′ ∘ 1_{𝔸} along F. Already (F′ ∘ G) ∘ F = F′ ∘ (G ∘ F) ≥ F′ ∘ 1_{𝔸} is obvious (using the unit of the adjunction); and if K′: 𝔹 [figure] 𝔹′ is another functor such that K′ ∘F ≥ F′ ∘1_{𝔸}, then K′ = K′ ∘1_{𝔹} ≥ K′ ∘(F ∘G) = (K′ ∘F)∘G ≥ (F′ ∘1_{𝔸})∘G = F′ ∘G. Thus, G is the absolute left Kan extension of 1_{𝔸} along F.

(2⇒3) being trivial, we now prove (3⇒1); it suffices to prove that F ⊣ ⟨1_{𝔸}, F ⟩. But the wanted unit ⟨1_{𝔸}, F ⟩∘F ≥ 1_{𝔸} is part of the universal property of the left Kan extension; and using the hypothesis that ⟨1_{𝔸}, F ⟩ is preserved by F itself, we have F ∘⟨1_{𝔸}, F ⟩ = ⟨F, F ⟩ which is smaller than 1_{𝔹} by the universal property of ⟨F, F ⟩ (since 1_{𝔹} ∘ F ≥ F).

The dual of this result says that a functor F: 𝔸 [figure] 𝔹 has a left adjoint in Cat(Q) if and only if the right Kan extension of 1_{𝔸} along F exists and is absolute, if and only if this right Kan extension exists and is preserved by F. Of course, the definition of right Kan extension is the dual of that of left Kan extension: given F: 𝔸 [figure] 𝔹 and G: 𝔸 [figure] ℂ in Cat(Q), the right Kan extension of F along G is a functor (F, G): ℂ [figure] 𝔹 such that (F, G)^{op}: ℂ^{op} [figure] 𝔹^{op} is the left Kan extension of F^{op} along G^{op} in Cat(Q^{op}). In elementary terms: there is a universal inequality (F, G) ∘ G ≤ F in Cat(Q).

After having introduced weighted colimits in a Q-category, we will discuss pointwise left Kan extensions: particular colimits that enjoy the universal property given above.

Skeletal categories

To any order corresponds an antisymmetric order by passing to equivalence classes of elements of the order. For Q-enriched categories this can be imitated: say that a Qcategory 𝔸 is skeletal if no two different objects in 𝔸 are isomorphic; equivalently this says that, for any Q-category ℂ, Cat(Q)(ℂ, 𝔸) is an antisymmetric order, i.e. that

Cat(Q)(ℂ, 𝔸) [figure] Dist(Q)(ℂ, 𝔸): F ↦ 𝔸(−, F −)

is injective. Any Q-category is then equivalent to its “skeletal quotient”.

4.7. Proposition. For a Q-category 𝔸, the following data define a skeletal Q-category 𝔸_{skel}:

- objects: (𝔸_{skel})_{0} = {[a] = {x ∈ 𝔸_{0} | x ≅ a}|a ∈ 𝔸_{0}}, with type function t[a] = ta;

- hom-arrows: for any [a], [a′] ∈ (𝔸_{skel})_{0}, 𝔸_{skel}([a′], [a]) = 𝔸(a′, a).

The object mapping [−]: 𝔸 [figure] 𝔸_{skel}: x ↦ [x] determines an equivalence in Cat(Q).

Proof. The construction of 𝔸_{skel} is well-defined, because:

- as a set the quotient of 𝔸_{0} by ≅ is well-defined (· ≅ · is an equivalence);

- t[a] is well-defined (all elements of the equivalence class [a] necessarily have the same type since they are isomorphic);

- for any a′ ≅ b′ and a ≅ b in 𝔸, 𝔸(b′, b) = 𝔸(a′, a), so 𝔸_{skel}([a′], [a]) is well-defined.

𝔸_{skel} inherits “composition” and “identities” from 𝔸, so it is a Q-category. And 𝔸_{skel} is skeletal: if [a] ≅ [a′] in (𝔸_{skel})_{0} then necessarily a ≅ a′ in 𝔸_{0}, which by definition implies that [a] = [a′].

The mapping 𝔸_{0} [figure] (𝔸_{skel})_{0}: x ↦ [x] determines a functor which is obviously fully faithful and essentially surjective on objects, so by 4.4 it is an equivalence in Cat(Q).

Skeletal Q-categories can be taken as objects of a full sub-2-category of Cat(Q), respectively a full sub-quantaloid of Dist(Q); the proposition above then says that the embedding Cat_{skel}(Q) [figure] Cat(Q) is a biequivalence of 2-categories, and Dist_{skel}(Q) [figure] Dist(Q) is then an equivalence of quantaloids. Cat_{skel}(Q) is locally antisymmetrically ordered, so the obvious 2-functor

Cat_{skel}(Q) [figure] Dist_{skel}(Q): (F: 𝔸 [figure] 𝔹) ↦ (𝔹(−, F −): 𝔸 [figure] 𝔹)

is the identity on objects, faithful (but not full), and locally fully faithful.

4.8. Example. The Q-categories PY and P^{†}X as defined in 3.9 are skeletal.

4.9. Example. Taking Q = 2, the skeletal 2-categories are precisely the antisymmetric orders, i.e. the partial orders (cf. 3.10).

4.10. Example. The Rel(Ω)-category associated to a sheaf on Ω, as in 3.11, is skeletal.

Later on we will encounter important Q-categories which are always skeletal: the categories P𝔸 and P^{†}𝔸 of (contravariant and covariant) presheaves on a given Q-category 𝔸, and also 𝔸’s Cauchy completion 𝔸_{cc}. (The reason, ultimately, that these Q-categories are skeletal, is that the hom-objects of the quantaloid Q are antisymmetrically ordered.)

5. Weighted (co)limits

Colimits

We consider a functor F: 𝔸 [figure] 𝔹 and a distributor Θ: ℂ [figure] 𝔸 between Q-categories. Since F determines a distributor 𝔹(F −, −): 𝔹 [figure] 𝔸, and by closedness of the quantaloid Dist(Q), we can calculate the universal lifting [Θ, 𝔹(F −, −)]: 𝔹 [figure] ℂ. A functor G: ℂ [figure] 𝔹 is the Θ-weighted colimit of F if it represents the universal lifting: 𝔹(G−, −) = [Θ, 𝔹(F −, −)]. If the Θ-weighted colimit of F exists, then it is necessarily essentially unique. It therefore makes sense to speak of “the” colimit and to denote it by colim(Θ, F); its universal property is thus that

𝔹( colim(Θ, F)−, −) = [Θ, 𝔹(F −, −)] in Dist(Q).

The following diagrams picture the situation:

[figure: 𝔸; F; 𝔹; ℂ; Θ; 𝔸; 𝔹; 𝔹(F −, −); [Θ, 𝔹(F −, −)]; ℂ; Θ; ≤; 𝔸; F; 𝔹; ℂ; Θ; colim(Θ, F)]

We wish to speak of “those Q-categories that admit all weighted colimits”, i.e. cocomplete Q-categories. But there is a small problem (the word is well-chosen).

5.1. Lemma. A cocomplete Q-category 𝔹 has at least as many objects as the base quantaloid Q.

Proof. For each X ∈ Q, consider the empty diagram in 𝔹 weighted by the empty distributor with domain ∗_{X}. Then colim(∅, ∅): ∗_{X} [figure] 𝔹 must exist by cocompleteness of 𝔹; it “picks out” an object of type X in 𝔹—thus 𝔹 must have such an object in the first place.

So, by the above, would the base quantaloid have a proper class of objects, then so would all cocomplete Q-enriched categories. This is a problem in the framework of this text, because we did not even bother defining “large” Q-categories, let alone develop a theory of Q-categories that is sensitive to these size-related issues. Therefore, from now on, we are happy to work with a small base quantaloid Q. The problem of “small” versus “large” then disappears: also “small” Q-categories can be cocomplete.

Next up is a collection of lemmas that will help us calculate colimits. (There will be some abuse of notation: for a distributor Φ: ∗_{A} [figure] 𝔸 and a functor F: 𝔸 [figure] 𝔹, colim(Φ, F) is in principle a functor from ∗_{A} to 𝔹. But such a functor simply “picks out” an object of type A in 𝔹. Therefore we will often think of colim(Φ, F) just as being that object. Of course, when the domain of the weight has more than one object, then any colimit with that weight is really a functor!)

5.2. Lemma.

1. For (Φ_{i}: ℂ [figure] 𝔸)_{i∈I} and F: 𝔸 [figure] 𝔹, if all colimits involved exist, then colim(⋁_{i} Φ_{i}, F) is the supremum of the colim(Φ_{i}, F) in the order Cat(Q)(ℂ, 𝔹):

colim(⋁_{i} Φ_{i}, F) ≅ ⋁_{i} colim(Φ_{i}, F).

2. For any F: 𝔸 [figure] 𝔹, colim(𝔸, F) (exists and) is isomorphic to F.

3. For Φ: 𝔻 [figure] ℂ, Θ: ℂ [figure] 𝔸 and F: 𝔸 [figure] 𝔹, suppose that the colimit colim(Θ, F) exists; then colim(Φ, colim(Θ, F)) exists if and only if colim(Θ ⊗_{ℂ} Φ, F) exists, in which case they are isomorphic.

4. For Θ: ℂ [figure] 𝔸 and F: 𝔸 [figure] 𝔹, colim(Θ, F) exists if and only if, for all objects c ∈ ℂ_{0}, colim(Θ(−, c), F) exists; then colim(Θ, F)(c) ≅ colim(Θ(−, c), F).

Proof. (1) By assumption,

ℂ( colim(⋁_{i} Φ_{i}, F)−, −) = [ ⋁_{i} Φ_{i}, 𝔹(F −, −)]
= ⋀_{i} [Φ_{i}, 𝔹(F −, −)]
= ⋀_{i} 𝔹( colim(Φ_{i}, F)−, −)

from which it follows that colim(⋁_{i} Φ_{i}, F) is the supremum of the colim(Φ_{i}, F) in the order Cat(Q)(ℂ, 𝔹).

(2) Trivially, [𝔸(−, −), 𝔹(F −, −)] = 𝔹(F −, −).

(3) By a simple calculation

[Φ, 𝔹( colim(Θ, F)−, −)] = [Φ, [Θ, 𝔹(F −, −)]] = [Θ ⊗_{ℂ} Φ, 𝔹(F −, −)]

so the Φ-weighted colimit of colim(Θ, F) and the Θ ⊗_{ℂ} Φ-weighted colimit of F are the same thing.

(4) Necessity is easy: colim(Θ, F): ℂ [figure] 𝔹 is a functor satisfying, for all c ∈ ℂ_{0} and b ∈ 𝔹_{0},

𝔹( colim(Θ, F)(c), b) = [Θ(−, c), 𝔹(F −, b)];

this literally says that colim(Θ, F)(c) is the Θ(−, c)-weighted colimit of F (which thus exists). As for sufficiency, we prove that the mapping

K: ℂ_{0} [figure] 𝔹_{0}: c ↦ colim(Θ(−, c), F)

is a functor: ℂ(c′, c) ≤ 𝔹(Kc′, Kc) because

𝔹(Kc′, Kc) = 𝔹(Kc′, −) ⊗_{𝔹} 𝔹(−, Kc)
= {𝔹(Kc, −), 𝔹(Kc′, −)}
= {[Φ(−, c), 𝔹(F −, −)], [Φ(−, c′), 𝔹(F −, −)]}

and, with slight abuse of notation^{4},

ℂ(c′, c) ≤ {[Φ(−, c), 𝔹(F −, −)], [Φ(−, c′), 𝔹(F −, −)]}
⟺ Φ(−, c′) ⊗_{∗_{tc′}} ℂ(c′, c) ⊗_{∗_{tc}} [Φ(−, c), 𝔹(F −, −)] ≤ 𝔹(F −, −)

4The Q-arrow ℂ(c′, c): tc [figure] tc′ is thought of as a one-element distributor ∗_{tc} [figure] ∗_{tc′}.

which holds because Φ(−, c′) ⊗_{∗_{t}c′} ℂ(c′, c) ≤ Φ(−, c) and by the universal property of the “[−, −]”. This functor K has, by construction, the required universal property for it to be the Φ-weighted colimit of F.

5.3. Lemma. For F ≤ G: 𝔸 [figure] 𝔹 and Θ: ℂ [figure] 𝔸, suppose that colim(Θ, F) and colim(Θ, G) exist; then colim(Θ, F) ≤ colim(Θ, G). In particular, suppose that F ≅ G; then colim(Θ, F) exists if and only if colim(Θ, G) exists, in which case they are isomorphic.

Proof. [−, −] is order-preserving in its second variable, so

𝔹(colim(Θ, F)−, −) = [Θ, 𝔹(F −, −)] ≥ [Θ, 𝔹(G−, −)] = 𝔹(colim(Θ, G)−, −),

hence colim(Θ, F) ≤ colim(Θ, G) as claimed. The second part of the statement is similar.

Let us point out two important consequences.

5.4. Corollary. A Q-category 𝔹 is cocomplete if and only if it admits colimits of the identity functor 1_{𝔹}: 𝔹 [figure] 𝔹 weighted by distributors into 𝔹 whose domains are one-object Q-categories^{5}: for every ϕ: ∗_{B} [figure] 𝔹 in Dist(Q), there exists an object b ∈ 𝔹, necessarily of type tb = B, such that

𝔹(b, −) = [ϕ, 𝔹] in Dist(Q).

Proof. One direction is trivial. For the other, consider a weight Θ: ℂ [figure] 𝔸 and a functor F: 𝔸 [figure] 𝔹. Then

[Θ, 𝔹(F −, −)] = [Θ, [𝔹(−, F −), 𝔹]] = [𝔹(−, F −) ⊗_{𝔸} Θ, 𝔹(1_{𝔹}−, −)]

so that colim(Θ, F): ℂ [figure] 𝔹 exists if and only if colim(𝔹(−, F −) ⊗_{𝔸} Θ, 1_{𝔹}): ℂ [figure] 𝔹 exists. But the latter exists if and only if, for all objects c ∈ ℂ,

colim (𝔹(−, F −) ⊗_{𝔸} Θ(−, c), 1_{𝔹}): ∗_{tc} [figure] 𝔹

exists—which they do by hypothesis.

5.5. Example. A 2-enriched category 𝔸 is an ordered set, and a distributor ϕ: ∗ [figure] 𝔸 is a down-closed subset of 𝔸. Since [ϕ, 𝔸] is the up-closed set of lower bounds of ϕ, it is representable if and only if the supremum of ϕ exists in 𝔸. Cocompleteness of 𝔸 as 2-category is thus the same thing as cocompleteness of 𝔸 as order (since it is equivalent to ask for the suprema of all subsets or only the suprema of down-closed sets).

5Later we will call those distributors contravariant presheaves, so 𝔹 is cocomplete if and only if it admits colimits of 1_{𝔹} weighted by contravariant presheaves on 𝔹.

5.6. Corollary. Given (F_{i}: 𝔸 [figure] 𝔹)_{i∈I} in Cat(Q), if the ⋁_{i} 𝔹(−, F_{i}−)-weighted colimit of 1_{𝔹} exists, then it is their supremum in the order Cat(Q)(𝔸, 𝔹):

⋁_{i} F_{i} ≅ colim ( ⋁_{i} 𝔹(−, F_{i}−), 1_{𝔹}).

Proof. The colimit, supposed to exist, satisfies

𝔹( colim(⋁_{i} 𝔹(−, F_{i}−), 1_{𝔹})−, −) = [ ⋁_{i} 𝔹(−, F_{i}−), 𝔹]
= ⋀_{i} [𝔹(−, F_{i}−), 𝔹]
= ⋀_{i} 𝔹(F_{i}−, −).

It follows straightforwardly that colim(⋁_{i} 𝔹(−, F_{i}−), 1_{𝔹}) is the supremum of the F_{i}.

This corollary thus says that, if there is a functor F: 𝔸 [figure] 𝔹 such that 𝔹(F −, −) = ⋀_{i} 𝔹(F_{i}−, −) for given (F_{i}: 𝔸 [figure] 𝔹)_{i∈I}, then F is the supremum of the F_{i}. In particular, for an object a ∈ 𝔸 it follows that F a is the supremum of the F_{i}a in the underlying order (𝔹, ≤).

5.7. Example. Concerning the result in 5.6, it is not true that each supremum in the order Cat(Q)(𝔸, 𝔹) is necessarily a weighted colimit in 𝔹. Consider the orders

[figure: 0; 1; a; b; c; d]

then the order morphisms f(0) = a, f(1) = c and g(0) = b, g(1) = c admit the supremum h(0) = h(1) = c, but clearly h(0) is not the supremum of f(0) and g(0). This implies that h is not the weighted colimit determined by f and g as in 5.6.

Limits

A limit in a Q-category 𝔸 is a colimit in the opposite category of 𝔸. This simple definition hides the difficulty that 𝔸 is a Q-category but 𝔸^{op} is a Q^{op}-category. Let us therefore make the definition of limit in 𝔸 elementary in terms of Q-enriched structures: for a distributor Φ: 𝔹 [figure] ℂ in Dist(Q) and a functor F: 𝔹 [figure] 𝔸 in Cat(Q), the Φ-weighted limit of F is – whenever it exists – a functor lim(Φ, F): ℂ [figure] 𝔸 in Cat(Q) such that

𝔸(−, lim(Φ, F) − ) = {Φ, 𝔸(−, F −)} in Dist(Q).

These diagrams picture the situation:

[figure: ℂ; 𝔹; Φ; F; 𝔸; ℂ; {Φ, 𝔸(−, F −)}; ℂ; Φ; 𝔸(−, F −); 𝔸; ≥; ℂ; lim(Φ, F); 𝔹; Φ; F; 𝔸]

A Q-category 𝔸 is complete when the Q^{op}-category 𝔸^{op} is cocomplete; by 5.4 we know that this is so precisely when for all ϕ: ∗_{A} [figure] 𝔸^{op} in Dist(Q^{op}), the ϕ-weighted colimit of 1_{𝔸^{op}} exists in Cat(Q^{op}); in terms of Q-enriched structures, this occurs precisely when for every^{6} ϕ: 𝔸 [figure] ∗_{A} in Dist(Q), there is an object a ∈ 𝔸, necessarily of type ta = A, such that

𝔸(−, a) = {ϕ, 𝔸} in Dist(Q).

5.8. Proposition. For any Q-category 𝔸 and any ϕ: 𝔸 [figure] ∗_{A} in Dist(Q), lim(ϕ, 1_{𝔸}) exists if and only if colim({ϕ, 𝔸}, 1_{𝔸}) exists, in which case they are isomorphic.

Proof. First suppose that a = lim(ϕ, 1_{𝔸}) exists: 𝔸(−, a) = {ϕ, 𝔸}. But then, using that 𝔸(−, a) ⊣ 𝔸(a, −) in Dist(Q), 𝔸(a, −) = [𝔸(−, a), 𝔸] = [{ϕ, 𝔸}, 𝔸], which says that a is the {ϕ, 𝔸}-weighted colimit of 1_{𝔸}.

Conversely, to say that a = colim({ϕ, 𝔸}, 1_{𝔸}) exists, means that 𝔸(a, −) = [{ϕ, 𝔸}, 𝔸]. Again using the adjunction 𝔸(−, a) ⊣ 𝔸(a, −), and by calculation with [−, −] and {−, −} in Dist(Q) it follows that 𝔸(−, a) = {[{ϕ, 𝔸}, 𝔸], 𝔸} = {ϕ, 𝔸}, so that indeed a is the ϕ-weighted colimit of 1_{𝔸}.

This proposition is valid for categories enriched in any quantaloid Q, hence also for Q^{op}- categories; translating the result in 5.8 for Q^{op}-categories in terms of Q-categories, we obtain its dual: for any ψ: ∗_{A} [figure] 𝔸 in Dist(Q), colim(ψ, 1_{𝔸}) exists if and only if lim([ψ, 𝔸], 1_{𝔸}) exists, in which case they are isomorphic.

5.9. Example. Applied to 2-enriched categories, i.e. ordered sets, 5.8 says that the infimum of an up-closed subset equals the supremum of the down-set of its lower bounds (and dually, the supremum of a down-closed subset equals the infimum of the up-closed set of its upper bounds).

The following is then an immediate consequence of 5.4 and 5.8.

5.10. Proposition. A Q-category is complete if and only if it is cocomplete.

This does not mean that “the theory of complete Q-categories” is the same as “the theory of cocomplete Q-categories” (although it is its dual): even though the objects are the same, the appropriate morphisms are not. For example, the free cocompletion of a Q-category is different from its free completion (see further).

6Later we will call such a distributor a covariant presheaf on 𝔸.

Absolute (co)limits and (co)continuous functors

Given a distributor Θ: 𝔸 [figure] 𝔹 and a functor F: 𝔹 [figure] ℂ between Q-categories for which colim(Θ, F): ℂ [figure] 𝔹 exists, a functor F′: 𝔹 [figure] 𝔹′ preserves this colimit if colim(Θ, F′ ∘ F) exists and is isomorphic to F′ ∘ colim(Θ, F). The functor F′ is cocontinuous if it preserves all colimits that exist in 𝔹. And colim(Θ, F) is absolute if it is preserved by any functor F′: 𝔹 [figure] 𝔹′.

Continuous functors and absolute limits are defined dually. For all of the following results we do not bother explicitly writing the dual statements, even though we may use them further on.

5.11. Proposition. If Θ: ℂ [figure] 𝔸 is a left adjoint in Dist(Q), then any colimit with weight Θ is absolute^{7}.

Proof. If Θ ⊣ Ψ in Dist(Q) and colim(Θ, F) exists, then necessarily

𝔹( colim(Θ, F)−, −) = [Θ, 𝔹(F −, −)] = Ψ ⊗_{𝔸} 𝔹(F −, −).

From this it is obvious that, for any F′: 𝔹 [figure] 𝔹′,

𝔹(F′ ∘ colim(Θ, F)−, −) = 𝔹( colim(Θ, F)−, −) ⊗_{𝔹} 𝔹′(F −, −)
= Ψ ⊗_{𝔸} 𝔹(F −, −) ⊗_{𝔹} 𝔹′(F′−, −)
= Ψ ⊗_{𝔸} 𝔹(F′ ∘ F −, −)
= [Θ, 𝔹(F′ ∘ F −, −)]

so colim(Θ, F′ ∘ F) exists and is isomorphic to F′ ∘ colim(Θ, F).

5.12. Proposition. If F′: 𝔹 [figure] 𝔹′ is a left adjoint in Cat(Q), then it is cocontinuous^{8}.

Proof. Suppose that F′ ⊣ G in Cat(Q)—i.e. 𝔹′(F′−, −) ⊣ 𝔹(G−, −) in Dist(Q). For any colimit

[figure: 𝔸; F; 𝔹; F′; 𝔹′; ℂ; Θ; colim(Θ, F)]

we can then calculate

𝔹′(F′ ∘ colim(Θ, F)−, −) = 𝔹( colim(Θ, F)−, −) ⊗_{𝔹} 𝔹′(F′−, −)

7A converse can be proved too: if any Θ-weighted colimit is absolute, then Θ is left adjoint [Street, 1983b].

8If 𝔹 is cocomplete, then also the converse holds—see 6.8.

= [Θ, 𝔹(F −, −)] ⊗_{𝔹} 𝔹′(F′−, −)
= {𝔹(G−, −), [Θ, 𝔹(F −, −)]}
= [Θ, {𝔹(G−, −), 𝔹(F −, −)}]
= [Θ, 𝔹(F −, −) ⊗_{𝔹} 𝔹′(F′−, −)]
= [Θ, 𝔹′(F′ ∘ F −, −)]

which shows that colim(Θ, F′ ∘ F) exists and equals F′ ∘ colim(Θ, F).

5.13. Corollary. Consider a fully faithful right adjoint G: 𝔸 [figure] 𝔹 in Cat(Q); if 𝔹 is cocomplete, then so is 𝔸.

Proof. Let us write F: 𝔹 [figure] 𝔸 for G’s left adjoint; now consider any colimit diagram

[figure: 𝔻; H; 𝔸; G; ⊥; 𝔹; F; ℂ; Φ; colim(Φ, G ∘ H)]

Then by cocontinuity of F we have F ∘ colim(Φ, G ∘ H) ≅ colim(Φ, F ∘ G ∘ H) (thus in particular the existence of the right hand side, the dotted arrow in the diagram), and by fully faithfulness of G (see 4.5), colim(Φ, F ∘ G ∘ H) ≅ colim(Φ, H).

The following two corollaries are now immediate.

5.14. Corollary. Consider a string of three adjoints in Cat(Q),

[figure: 𝔸; F; H; ⊥; ⊥; 𝔹; G] ,

with F fully faithful. If 𝔹 is cocomplete, then so is 𝔸.

5.15. Corollary. If F: 𝔸 [figure] 𝔹 is an equivalence in Cat(Q), then both F and its inverse equivalence are cocontinuous, and 𝔸 is cocomplete if and only if 𝔹 is.

6. Presheaves

Contravariant presheaves

Principally the “contravariant presheaves on a Q-category 𝔸” should be the objects of a Q-category P𝔸 which “classifies” distributors into 𝔸:

for every Q-category ℂ, Dist(Q)(ℂ, 𝔸) ≃ Cat(Q)(ℂ, P𝔸).  (11)

(Here we ask a priori for an equivalence of orders, but it will prove to be an isomorphism of sup-lattices.) Putting ℂ = ∗_{C} in the above equivalence dictates that a contravariant presheaf of type C ∈ Q on 𝔸 is a distributor ∗_{C} [figure] 𝔸.

In what follows we will systematically use small Greek letters – instead of capital ones – for such contravariant presheaves; the elements of such a ϕ: ∗_{C} [figure] 𝔸 will be written as ϕ(a): C [figure] ta (instead of ϕ(a, ∗), thus stressing the fact that the action of ∗_{C} on ϕ is trivial).

The precise sense in which a distributor ϕ: ∗_{C} [figure] 𝔸 is contravariant, is the following: the action of 𝔸 on ϕ can be written as 𝔸(a′, a) ≤ {ϕ(a), ϕ(a′)} (for all a, a′ ∈ 𝔸_{0}). A contravariant presheaf on 𝔸 may thus be thought of as a “functor on 𝔸^{op} with values in Q” sending a to ϕ(a). To make this precise: indeed, a ↦ ϕ(a) is a functor between 𝔸^{op} and (P^{†}C)^{op} (see 3.9) in Cat(Q^{op}), or equivalently a functor between 𝔸 and P^{†}C in Cat(Q)^{co} (by the duality in 3.8).

Since we agreed to work over a small base quantaloid Q, the collection of contravariant presheaves on a Q-category 𝔸 forms a Q-typed set^{9}. We now need to enrich this set in Q in such a way that the classifying property in (11) holds.

6.1. Proposition. For any Q-category 𝔸, the following defines the Q-category P𝔸 of contravariant presheaves on 𝔸:

- objects: (P𝔸)_{0} = {ϕ: ∗_{C} [figure] 𝔸 in Dist(Q) | C ∈ Q} with types t(ϕ: ∗_{C} [figure] 𝔸) = C;

- hom-arrows: for ϕ: ∗_{C} [figure] 𝔸 and ψ: ∗_{C′} [figure] 𝔸 put P𝔸(ψ, ϕ): C [figure] C′ to be the single

element of the distributor [ψ, ϕ]: ∗_{C} [figure] ∗_{C′}.

This Q-category is skeletal, and it satisfies the equivalence in (11) as follows:

- for a distributor Φ: ℂ [figure] 𝔸 the corresponding functor F_{Φ}: ℂ [figure] P𝔸 maps c ∈ ℂ_{0} onto Φ(−, c): tc [figure] 𝔸;

- and for a functor F: ℂ [figure] P𝔸 the corresponding distributor Φ_{F}: ℂ [figure] 𝔸 has elements, for (c, a) ∈ ℂ_{0} × 𝔸_{0}, Φ_{F} (a, c) = F (c)(a).

Proof. With the properties of “[−, −]” in Dist(Q) it is easily verified that P𝔸 is a Q-category. For ϕ, ψ ∈ P𝔸, observe that ϕ ≤ ψ in (P𝔸)_{0} if and only if tϕ = tψ =: C and (1_{C}) ≤ P𝔸(ϕ, ψ), that is to say, ϕ, ψ: ∗_{C} [figure] 𝔸 and 1_{C} ≤ [ϕ, ψ] in Dist(Q), thus precisely when ϕ ≤ ψ: ∗_{C} [figure] 𝔸. This shows that the underlying order on the objects of P𝔸 (which are certain distributors) coincides with their order as arrows in Dist(Q)—so it is antisymmetric, i.e. P𝔸 is a skeletal Q-category. As for the equivalence of orders in (11), it is easily seen that the mappings F ↦ Φ_{F} and Φ ↦ F_{Φ} as in the statement of the proposition define a bijection:

9Would the objects of the base category Q form a class, then the collection of contravariant presheaves on 𝔸 would be large! The reason – ultimately – is that P𝔸 is a cocomplete category—see 5.1 and 6.4.

- for any c ∈ ℂ_{0}, F_{Φ_{F}} (c) = Φ_{F} (−, c) = F (c), so F_{Φ_{F}} = F,
- for any (c, a) ∈ ℂ_{0} × 𝔸_{0}, Φ_{F_{Φ}}(a, c) = F_{Φ}(c)(a) = Φ(a, c), so Φ_{F_{Φ}} = Φ.

This bijection preserves order in both directions:

Φ ≤ Ψ in Dist(ℂ, 𝔸)

⟺ ∀(c, a) ∈ ℂ_{0} × 𝔸_{0} : Φ(a, c) ≤ Ψ(a, c) in Q(tc, ta) ⟺ ∀c ∈ ℂ_{0} : F_{Φ}(c) ≤ F_{Ψ}(c) in ((P𝔸)_{0}, ≤) ⟺ F_{Φ} ≤ F_{Ψ} in Cat(Q)(ℂ, P𝔸)

and likewise

F ≤ G in Cat(Q)(ℂ, P𝔸)
⟺ ∀c ∈ ℂ_{0} : F (c) ≤ G(c) in ((P𝔸)_{0}, ≤)
⟺ ∀(c, a) ∈ ℂ_{0} × 𝔸_{0} : Φ_{F} (a, c) ≤ Φ_{G}(a, c) in Q(tc, ta)
⟺ Φ_{F} ≤ Φ_{G} in Dist(ℂ, 𝔸).

6.2. Example. For a Q-object Y, the Q-category of contravariant presheaves on ∗_{Y} is denoted by PY instead of P(∗_{Y})—see also 3.9.

Very often there will be an abuse of notation: for ϕ: ∗_{C} [figure] 𝔸 and ψ: ∗_{C′} [figure] 𝔸 there is a distributor [ψ, ϕ]: ∗_{C} [figure] ∗_{C′} with precisely one element (because it goes between one-object categories); therefore we defined P𝔸(ψ, ϕ) to be the single element of this distributor, for we want P𝔸(ψ, ϕ) to be a Q-arrow. But more often than not we will simply write P𝔸(ψ, ϕ) = [ψ, ϕ], thus identifying a quantaloid Q with the full subcategory of Dist(Q) determined by the one-object Q-categories whose hom-arrows are identities (see also (9) and the appendix).

The equivalence of orders in (11) has a nice side-effect: because P𝔸 is skeletal, Cat(Q)(ℂ, P𝔸) is an antisymmetric order which is order-isomorphic to the sup-lattice Dist(Q)(ℂ, 𝔸), thus it is a sup-lattice too. In particular are F ↦ Φ_{F} and Φ ↦ F_{Φ} sup-morphisms.

To the identity distributor 𝔸: 𝔸 [figure] 𝔸 corresponds, under the equivalence in (11), the functor

Y_{𝔸}: 𝔸 [figure] P𝔸 : a ↦ (𝔸(−, a): ∗_{ta} [figure] 𝔸)

which is, of course, the Yoneda embedding for 𝔸. The presheaf 𝔸(−, a): ∗_{ta} [figure] 𝔸 is said to be represented by a ∈ 𝔸_{0}. We can now prove the “Yoneda lemma” for presheaves on a Q-enriched category.

6.3. Proposition. For any Q-category 𝔸, any object a ∈ 𝔸_{0} and any ϕ: ∗_{C} [figure] 𝔸 in Dist(Q), P𝔸(Y_{𝔸}a, ϕ) = ϕ(a). As a consequence, the Yoneda embedding Y_{𝔸}: 𝔸 [figure] P𝔸 is fully faithful.

Proof. Using that 𝔸(−, a) ⊣ 𝔸(a, −) in Dist(Q), we have by definition of P𝔸(−, −) that P𝔸(Y_{𝔸}a, ϕ) = [𝔸(−, a), ϕ] = 𝔸(a, −) ⊗_{𝔸} ϕ = ϕ(a).

An – almost – immediate but non the less important consequence of 6.1 is the following.

6.4. Proposition. The category P𝔹 of contravariant presheaves on a Q-category 𝔹 is cocomplete.

Proof. Given a distributor Θ: ℂ [figure] 𝔸 and a functor F: 𝔸 [figure] P𝔹, to F corresponds a unique distributor Φ_{F}: 𝔸 [figure] 𝔹 and, in turn, to Φ_{F} ⊗_{𝔸} Θ: ℂ [figure] 𝔹 corresponds a unique functor F_{Φ_{F} ⊗_{𝔸}Θ}: ℂ [figure] P𝔹. We claim that this latter functor is the colimit of F weighted by Θ: for any c ∈ ℂ and β ∈ P𝔹 we should verify that

P𝔹(F_{Φ_{F} ⊗_{𝔸}Θ}(c), β) = [Θ(−, c), P𝔹(F −, β)].

The right hand side equals [Θ(−, c), [F −, β]] and the left hand side equals

[(Φ_{F} ⊗_{𝔸} Θ)(−, c), β] = [Φ_{F} ⊗_{𝔸} Θ(−, c), β] = [Θ(−, c), [Φ_{F} , β]].

Since [F −, β] = [Φ_{F} , β], the required universality holds.

6.5. Corollary. For any ϕ: ∗_{C} [figure] 𝔸, the ϕ-weighted colimit of Y_{𝔸}: 𝔸 [figure] P𝔸 (exists and) is ϕ itself.

In other words, every contravariant presheaf on a Q-category 𝔸 is canonically the colimit of representable presheaves: colim(ϕ, Y_{𝔸}) = ϕ (and we may write an equality instead of merely an isomorphism, for P𝔸 is skeletal).

Here is another result that should sound familiar.

6.6. Proposition. The Yoneda embedding Y_{𝔸}: 𝔸 [figure] P𝔸 is continuous.

Proof. Consider a weighted limit (which happens to exist) in 𝔸, like so:

[figure: ℂ; lim(Φ, F); 𝔹; Φ; F; 𝔸; Y_{𝔸}; P𝔸]

Then, for any c ∈ ℂ_{0},

(Y_{𝔸} ∘ lim(Φ, F))(c) = 𝔸(−, lim(Φ, F)(c)) = {Φ(c, −), 𝔸(−, F −)}

and therefore

P𝔸(−, Y_{𝔸} ∘ lim(Φ, F)−) = [−, {Φ, 𝔸(−, F −)}]
= {Φ, [−, 𝔸(−, F −)]}
= {Φ, P𝔸(−, Y_{𝔸} ∘ F −)}

so Y_{𝔸} ∘ lim(Φ, F) is lim(Φ, Y_{𝔸} ∘ F).

When Y_{𝔸}: 𝔸 [figure] P𝔸 is a right adjoint, then 𝔸 is cocomplete (combine 5.13 with 6.3 and 6.4). Also the converse can be shown; actually, P𝔸 is the free cocompletion of 𝔸 in Cat(Q). To make this precise in an elegant way, it is useful to treat pointwise left Kan extensions first.

Pointwise left Kan extensions

If for F: 𝔹 [figure] 𝔸 and G: 𝔹 [figure] ℂ in Cat(Q) the ℂ(G−, −)-weighted colimit of F exist, then it is the left Kan extension of F along G; so ⟨F, G⟩ ≅ colim(ℂ(G−, −), F) whenever the right hand side exists. By lemma 5.2 we know that in this case the action of ⟨F, G⟩: ℂ [figure] 𝔹 on objects is given by the colimit formula ⟨F, G⟩(c) ≅ colim(ℂ(G−, c), F). Such a colimit weighted by the right adjoint distributor determined by some functor is therefore said to be a pointwise left Kan extension. (Note that, for a pointwise left Kan extension, the two notions of absoluteness, one relevant for Kan extensions and the other relevant for colimits, coincide.)

A Q-category is said to admit all pointwise left Kan extensions when it admits all colimits of this special kind. A cocomplete Q-category then certainly admits all pointwise left Kan extensions. (For a cocomplete 𝔹, any functor G: 𝔸 [figure] ℂ thus determines an adjunction

⟨−, G⟩ ⊣ (− ∘ G): Cat(Q)(ℂ, 𝔹) [figure] Cat(Q)(𝔸, 𝔹).

This has been called the “meta-adjointness” associated to the concept of Kan extension.) But it is also true that a Q-category admitting all pointwise Kan extensions, is cocomplete! This follows from 5.4 in combination with 6.9 below; note that the Yoneda embedding plays a crucial rôle.

6.7. Proposition. Suppose that for F: 𝔹 [figure] 𝔸 and G: 𝔹 [figure] ℂ in Cat(Q) the pointwise left Kan extension of F along G exists. If G is fully faithful, then ⟨F, G⟩ ∘ G ≅ F.

Proof. It is easily seen that, ⟨F, G⟩ being a pointwise left Kan extension,

⟨F, G⟩ ∘ G ≅ F
⟺ 𝔹(⟨F, G⟩ ∘ G−, −) = 𝔹(F −, −)
⟺ ℂ(G−, −) ⊗_{ℂ} 𝔹(⟨F, G⟩−, −) = 𝔹(F −, −)
⟺ ℂ(G−, −) ⊗_{ℂ} [ℂ(G−, −), 𝔹(F −, −)] = 𝔹(F −, −)
⟺ ℂ(G−, −) ⊗_{ℂ} ℂ(−, G−) ⊗_{𝔸} 𝔹(F −, −) = 𝔹(F −, −)
⟺ ℂ(G−, G−) ⊗_{𝔸} 𝔹(F −, −) = 𝔹(F −, −)

so if G is fully faithful, i.e. ℂ(G−, G−) = 𝔸, then the above holds.

6.8. Proposition. Let 𝔸 be a cocomplete Q-category; then the cocontinuous functors with domain 𝔸 coincide with the left adjoints.

Proof. One direction is a consequence of 5.12. To prove the other, let F: 𝔸 [figure] 𝔹 be a cocontinuous functor. By cocompleteness of 𝔸, ⟨1_{𝔸}, F ⟩ not only exists but is even pointwise, i.e. it is a colimit, so it must be preserved by F. By 4.6 we conclude that F is a left adjoint.

6.9. Proposition. A functor F: 𝔸 [figure] 𝔹 has a pointwise left Kan extension along Y_{𝔸}: 𝔸 [figure] P𝔸 if and only if, for every ϕ ∈ P𝔸, the ϕ-weighted colimit of F exists. In this case, ⟨F, Y_{𝔸}⟩: P𝔸 [figure] 𝔹 is left adjoint to ⟨Y_{𝔸}, F ⟩: 𝔹 [figure] P𝔸, and ⟨F, Y_{𝔸}⟩∘Y_{𝔸} ≅ F. Actually, ⟨F, Y_{𝔸}⟩ is then the essentially unique cocontinuous factorization of F through Y_{𝔸}.

Proof. If by hypothesis, ⟨F, Y_{𝔸}⟩ = colim(P𝔸(Y_{𝔸}−, −), F) exists, then ⟨F, Y_{𝔸}⟩(ϕ) = colim(P𝔸(Y_{𝔸}−, ϕ), F) = colim(ϕ, F) because P𝔸(Y_{𝔸}−, ϕ) = ϕ. Conversely, supposing that 𝔹 allows colimits of F weighted by presheaves on 𝔸, colim(P𝔸(Y_{𝔸}−, −), F) exists because, for all ϕ = P𝔸(Y_{𝔸}−, ϕ), colim(P𝔸(Y_{𝔸}−, ϕ), F) exists (see 5.2).

Since P𝔸 is cocomplete, ⟨Y_{𝔸}, F ⟩ exists, is pointwise, and can thus be computed as ⟨Y_{𝔸}, F ⟩(b) ≅ colim(𝔹(F −, b), Y_{𝔸}) ≅ 𝔹(F −, b). It is then a matter of applying 4.2 to prove the adjunction:

⟨F, Y_{𝔸}⟩ ⊣ ⟨Y_{𝔸}, F ⟩
⟺ ∀ϕ ∈ P𝔸, ∀b ∈ 𝔹 : 𝔹(⟨F, Y_{𝔸}⟩(ϕ), b) = P𝔸(ϕ, ⟨Y_{𝔸}, F ⟩(b))
⟺ ∀ϕ ∈ P𝔸, ∀b ∈ 𝔹 : 𝔹( colim(ϕ, F), b) = [ϕ, 𝔹(F −, b)]

which holds by the definition of colimit.

Finally, by fully faithfulness of Y_{𝔸} and 6.7 it follows that F factors through Y_{𝔸} as ⟨F, Y_{𝔸}⟩. Would now G: P𝔸 [figure] 𝔹 be another cocontinuous functor such that F ≅ G ∘ Y_{𝔸} then – by cocontinuity and using the fact that any presheaf ϕ ∈ P𝔸 is canonically the colimit of representables – G and ⟨F, Y_{𝔸}⟩ agree on all objects of P𝔸:

G(ϕ) = G(colim(ϕ, Y_{𝔸}))
≅ colim(ϕ, G ∘ Y_{𝔸})
≅ colim(ϕ, F)
= ⟨F, Y_{𝔸}⟩(ϕ).

6.10. Corollary. A Q-category 𝔸 is cocomplete if and only if Y_{𝔸}: 𝔸 [figure] P𝔸 is a right adjoint in Cat(Q).

Proof. If Y_{𝔸} is right adjoint, apply 5.13 to Y_{𝔸}, using 6.3 and 6.4. If 𝔸 is cocomplete, apply 6.9 to the identity functor 1_{𝔸}: 𝔸 [figure] 𝔸: ⟨1_{𝔸}, Y_{𝔸}⟩ ⊣ ⟨Y_{𝔸}, 1_{𝔸}⟩ = Y_{𝔸}.

Free cocompletion

Applying 6.9 we easily obtain that, for two Q-categories 𝔸 and 𝔹, with 𝔹 cocomplete, any functor F: 𝔸 [figure] 𝔹 factors through Y_{𝔸}: 𝔸 [figure] P𝔸 by the essentially unique cocontinuous – even left adjoint – functor ⟨F, Y_{𝔸}⟩: P𝔸 [figure] 𝔹. Write now Cocont(Q) for the sub-2-category of Cat(Q) whose objects are the cocomplete Q-categories, and whose morphisms are the cocontinuous functors, or equivalently (by 6.8) the left adjoint ones; there is a forgetful 2-functor U: Cocont(Q) [figure] Cat(Q). The above then says that, for Q-categories 𝔸 and 𝔹, with 𝔹 cocomplete, the Yoneda embedding Y_{𝔸}: 𝔸 [figure] P𝔸 induces a natural equivalence of orders

− ∘ Y_{𝔸}: Cocont(Q)(P𝔸, 𝔹) [figure: ∼] Cat(Q)(𝔸, 𝔹)

with inverse F ↦ ⟨F, Y_{𝔸}⟩.

6.11. Proposition. The presheaf construction provides a left (bi)adjoint to the inclusion of Cocont(Q) in Cat(Q): P ⊣ U: Cocont(Q) [figure] Cat(Q).

For any two Q-categories 𝔸 and 𝔹, we have – by 6.1 and the above – equivalences

Dist(Q)(𝔸, 𝔹) [figure: ∼] Cat(Q)(𝔸, P𝔹) [figure: ∼] Cocont(Q)(P𝔸, P𝔹)  (12)

sending a Φ: 𝔸 [figure] 𝔹 to ⟨F_{Φ}, Y_{𝔸}⟩: P𝔹 [figure] P𝔸 (with F_{Φ}: 𝔹 [figure] P𝔸: b ↦ Φ(−, b) as in 6.1). We can prove more.

6.12. Proposition. The locally ordered^{10} category Cocont(Q) has stable local colimits, and the action

Dist(Q) [figure] Cocont(Q): (Φ: 𝔸 [figure] 𝔹) ↦ (⟨F_{Φ}, Y_{𝔸}⟩: P𝔸 [figure] P𝔹)

is a 2-functor which is locally an equivalence.

Proof. We already know that Cocont(Q) is a 2-category; we must show that it has stable local colimits. By 5.6 we know how to compute suprema in Cat(Q)(𝔸, 𝔹) whenever 𝔹 is cocomplete; we may thus apply this to given arrows (F_{i}: 𝔸 [figure] 𝔹)_{i∈I} in Cocont(Q):

⋁_{i∈I} F_{i} := colim ( ⋁_{i∈I} 𝔹(−, F_{i}−), 1_{𝔹}).

But we must show that ⋁_{i} F_{i} is still a cocontinuous functor (for a priori it is merely a functor); equivalently (by cocompleteness of 𝔸), we may assume that each F_{i} is a left adjoint in Cat(Q), and show that ⋁_{i} F_{i} itself is a left adjoint in Cat(Q). Let us write F_{i} ⊣ G_{i} for such adjunctions; then, by 5.6 and its dual^{11},

𝔹(⋁_{i} F_{i}−, −) = ⋀_{i} 𝔹(F_{i}−, −) = ⋀_{i} 𝔸(−, G_{i}−) = 𝔸(−, ⋀_{i} G_{i}−).

10... but not antisymmetrically so!

11The dual of 5.6 says the following: if 𝔸 is a complete category, then the infimum of functors (G_{i}: 𝔹 [figure] 𝔸)_{i∈I} exists in Cat(Q)(𝔹, 𝔸), and is given by the ⋁_{i} 𝔸(G_{i}−, −)-weighted limit of 1_{𝔸}. This means exactly that this infimum ⋀_{i} G_{i} satisfies 𝔸(−, ⋀_{i} G_{i}−) = ⋀_{i} 𝔸(−, G_{i}−). By 5.10 any complete category is also cocomplete, so this applies in this case!

That is to say, ⋁_{i} F_{i} ⊣ ⋀_{i} G_{i} in Cat(Q). As for the stability of these local colimits, take for example (F_{i}: 𝔸 [figure] 𝔹)_{i∈I} and G: 𝔹 [figure] ℂ in Cocont(Q); then

G ∘ (⋁_{i} F_{i}) = G ∘ colim ( ⋁_{i} 𝔹(−, F_{i}−), 1_{𝔹})
≅ colim ( ⋁_{i} 𝔹(−, F_{i}−), G)
≅ colim (ℂ(−, G−) ⊗_{𝔹} ⋁_{i}(−, F_{i}−), 1_{ℂ})
≅ colim ( ⋁_{i}(−, G ∘ F_{i}−), 1_{ℂ})
= ⋁_{i}(G ∘ F_{i})

because G is cocontinuous. Now consider F: 𝔸 [figure] 𝔹 and (G_{i}: 𝔹 [figure] ℂ)_{i∈I}, then for any a ∈ 𝔸_{0}

((⋁_{i} G_{i}) ∘ F )(a) = colim ( ⋁_{i} ℂ(−, G_{i}−), 1_{ℂ})(F a)
≅ colim ( ⋁_{i} ℂ(−, G_{i}−) ⊗_{𝔹} 𝔹(−, F a), 1_{ℂ})
≅ colim ( ⋁_{i} ℂ(−, G_{i}(F a)), 1_{ℂ})
≅ colim ( ⋁_{i} ℂ(−, G_{i} ∘ F −), 1_{ℂ})(a)
= ( ⋁_{i}(G_{i} ∘ F))(a)

by properties of the colimit – see 5.2 – so (⋁_{i} G_{i}) ∘ F = ⋁_{i}(G_{i} ∘ F).

To prove the functoriality of the action described in the statement of the proposition, observe that for Φ: 𝔸 [figure] 𝔹 and θ ∈ P𝔸,

⟨F_{Φ}, Y_{𝔸}⟩(θ) = colim (P𝔸(θ, Y_{𝔸}−), F_{Φ})
= colim(θ, F_{Φ})
= Φ ⊗_{𝔸} θ.

In the first line we use the cocompleteness of P𝔹 to compute the (pointwise) left Kan extension as a colimit; to pass from the first line to the second we use the Yoneda lemma 6.3; the third line follows from the way in which weighted colimits are calculated in a presheaf category, cf. the proof of 6.4. Using this formula it is clear that Φ ↦ ⟨F_{Φ}, Y_{𝔸}⟩ is functorial. That this functor is really a 2-functor, and that it is locally an equivalence, follows from the preceding remarks, in particular (12).

Restricting to skeletal cocomplete Q-categories, Cocont_{skel}(Q) is a quantaloid and the 2-functor above corestricts to a quantaloid homomorphism Dist(Q) [figure] Cocont_{skel}(Q) that induces sup-isomorphisms between the hom-sup-lattices.

6.13. Example. For a 2-category 𝔸, i.e. an ordered set, P𝔸 is set of down-sets of 𝔸, ordered by inclusion. The Yoneda embedding Y_{𝔸}: 𝔸 [figure] P𝔸 sends an element to the principal down-set that it determines. An eventual left adjoint to Y_{𝔸} is “taking the supremum”. The isomorphism Dist(2)(𝔸, 𝔹) ≅ Cocont_{skel}(2)(P𝔸, P𝔹) says that an ideal relation between given orders is the same thing as a sup-morphism between the sup-lattices of down-sets of those orders. In fact, Cocont_{skel}(2) = Sup.

Covariant presheaves, free completion

The results in the previous sections may be dualized; one obtains that “the Q-category of covariant presheaves on 𝔸 is 𝔸’s free completion in Cat(Q)”. Let us quickly do this exercise.

A covariant presheaf of type C ∈ Q on a Q-category 𝔸 is a distributor ϕ: 𝔸 [figure] ∗_{C}. Covariant presheaves form the object-set of a Q-category P^{†}𝔸, whose hom-arrows are, for ϕ: 𝔸 [figure] ∗_{C} and ψ: 𝔸 [figure] ∗_{C′},

P^{†}𝔸(ψ, ϕ) = single element of {ϕ, ψ} in Dist(Q).

This Q-category satisfies

for every Q-category ℂ, Dist(Q)(𝔸, ℂ) ≃ Cat(Q)^{co}(ℂ, P^{†}𝔸),  (13)

and it is complete. There is a Yoneda embedding Y_{𝔸}^{†}: 𝔸 [figure] P^{†}𝔸, sending an object a ∈ 𝔸 to the representable distributor 𝔸(a, −): 𝔸 [figure] ∗_{ta}, which is fully faithful. 𝔸 is complete if and only if Y_{𝔸}^{†} is a left adjoint in Cat(Q); but in any case is Y_{𝔸}^{†} cocontinuous.

Continuous functors whose domain is a complete Q-category, coincide with the right adjoints. Denoting Cont(Q) for the sub-2-category of Cat(Q) whose objects are complete Q-categories and arrows are continuous functors, we have

P^{†} ⊣ U: Cont(Q) [figure] Cat(Q);

that is, P^{†}𝔸 is the free completion of 𝔸 in Cat(Q). Furthermore, Cont(Q) has stable local limits, and there is a 2-functor Dist(Q) [figure] Cont(Q)^{coop}, which is locally an equivalence, sending a distributor Φ: 𝔸 [figure] 𝔹 to the right Kan extension (F^{Φ}, Y_{𝔸}^{†}): P^{†}𝔹 [figure] P^{†}𝔸, where F^{Φ}: 𝔹 [figure] P^{†}𝔸 is determined by Φ through the equivalence in (13).

Since a Q-category is complete if and only if it is cocomplete, and since between such categories continuous (cocontinuous) functors coincide with right adjoints (left adjoints), there is an isomorphism of (locally cocompletely ordered) 2-categories

Cocont(Q) [figure: ∼] Cont(Q)^{coop}: (F: 𝔸 [figure] 𝔹) [figure] (F^{∗}: 𝔹 [figure] 𝔸)

where F ⊣ F^{∗} in Cat(Q). Considering only skeletal objects, this restricts to an isomorphism of quantaloids Cocont_{skel}(Q) ≅ Cont_{skel}(Q)^{coop}.

Assembling previous results we obtain that: for a Q-category 𝔸, Y_{𝔸}: 𝔸 [figure] P𝔸 is a right adjoint if and only if Y_{𝔸}^{†}: 𝔸 [figure] P^{†}𝔸 is a left adjoint; this is precisely the case when 𝔸 is (co)complete. It is interesting to take a closer look at this result. First observe that, in any case, there is an adjunction

\xymatrix{ & [−, 𝔸] \ar[dr] \\ P𝔸 & ⊥ & P^{†}𝔸 \\  & \{−, 𝔸\} \ar[ul]}

in Cat(Q), where the functors involved are obtained by calculating {−, −} and [−, −] in Dist(Q):

[−, 𝔸]: P𝔸 [figure] P^{†}𝔸: (ϕ: ∗_{A} [figure] 𝔸) ↦ ([ϕ, 𝔸]: 𝔸 [figure]∗_{A} ),
{−, 𝔸}: P^{†}𝔸 [figure] P𝔸: (ψ: 𝔸 [figure]∗_{A} ) ↦ ({ψ, 𝔸}: ∗_{A} [figure] 𝔸).

Now considering the diagram of Q-categories and functors

[figure: 𝔸; Y_{𝔸}^{†}; Y_{𝔸}; P^{†}𝔸; {−, 𝔸}; R; P𝔸; [−, 𝔸]; L]

we have on the one hand that, if L ⊣ Y_{𝔸} is known, then R := L ∘ {−, 𝔸} is right adjoint to Y_{𝔸}^{†}; and dually, if Y_{𝔸}^{†} ⊣ R is known, then L := R ∘ [−, 𝔸] is left adjoint to Y_{𝔸}. Actually, L(ϕ) = colim(ϕ, 1_{𝔸}) and R(ψ) = lim(ψ, 1_{𝔸}) (whenever one, thus also the other, exists), so that the crux of this construction is really the statement in 5.8.

6.14. Example. Take 2-categories, then Cont(2) ≃ Inf = Cont_{skel}(2). For a 2-category 𝔸, P^{†}𝔸 is the set of up-sets of 𝔸, ordered by containment (reverse inclusion); Y_{𝔸}^{†} sends an element of 𝔸 to the corresponding principal up-set. An eventual right adjoint to Y_{𝔸}^{†} is “infimum”. It is well-known that Sup ≅ Inf^{coop} in QUANT by passing from left-adjoints to right-adjoints in Cat_{skel}(2) = Ord. The adjunction between P𝔸 (down-sets of 𝔸, ordered by inclusion) and P^{†}𝔸 (up-sets of 𝔸, ordered by containment) is constituted by the left adjoint that sends a down-set to the up-set of its upper bounds, and the right adjoint that sends an up-set to the down-set of its lower bounds.

6.15. Example. For a Q-object X, the category of covariant presheaves on ∗_{X} is denoted by P^{†}X rather than P^{†}(∗_{X}) (cf. 3.9).

7. Cauchy completion and Morita equivalence

Cauchy complete categories

A left adjoint distributor Φ: ℂ [figure] 𝔸 is often called a Cauchy distributor (into 𝔸); we will systematically denote its right adjoint by Φ^{∗}:

[figure: 𝔸; ⊥; Φ; 𝔹; Φ^{∗}] .

In particular, a Cauchy presheaf ϕ: ∗_{C} [figure] 𝔸 is a contravariant presheaf with a right adjoint ϕ^{∗}: 𝔸 [figure] ∗_{C}, whose elements are ϕ^{∗}(a): ta [figure] C.

A Cauchy distributor Φ: ℂ [figure] 𝔸 is said to converge (to a functor F: ℂ [figure] 𝔸) if (there exists a functor F: ℂ [figure] 𝔸 such that) Φ = 𝔹(−, F −)—or equivalently, Φ^{∗} = 𝔹(F −, −). A Cauchy distributor Φ: ℂ [figure] 𝔸 need not converge, but if it does then the functor F: ℂ [figure] 𝔸 that it converges to, is essentially unique. We are interested in those Q-categories in which “all Cauchy distributors converge”: say that 𝔸 is Cauchy complete if, for all ℂ, any Cauchy distributor Φ: ℂ [figure] 𝔸 converges. In other words, letting Map(Dist(Q))(ℂ, 𝔸) stand for the full subcategory of Dist(Q)(ℂ, 𝔸) whose objects are the Cauchy distributors (“maps”), 𝔸 is Cauchy complete if and only if, for every ℂ, the mapping

Cat(Q)(ℂ, 𝔸) [figure] Map(Dist(Q))(ℂ, 𝔸): F ↦ 𝔸(−, F −)

is surjective; recalling that it is always “essentially injective”, we thus ask for it to be an equivalence.

Fortunately, to have Cauchy completeness, convergence of the Cauchy presheaves suffices!

7.1. Proposition. A Q-category 𝔸 is Cauchy complete if and only if all Cauchy presheaves on 𝔸 converge (i.e. are representable).

Proof. We only need to prove the “if”. Let Φ ⊣ Φ^{∗}: 𝔸 [figure] ℂ, then for each c ∈ ℂ_{0}, Φ(−, c) ⊣ Φ^{∗}(c, −): 𝔸 [figure] ℂ, because Φ⊗_{ℂ}ℂ(−, c) = Φ(−, c) and ℂ(c, −)⊗_{ℂ}Φ^{∗} = Φ^{∗}(c, −), and by composition of adjoints. By assumption there exists for each c ∈ ℂ_{0} a functor F c: ∗_{tc} [figure] 𝔸 – that we will henceforth identify with the single element of type tc that it picks out – such that 𝔸(−, F c) = Φ(−, c) or equivalently, 𝔸(F c, −) = Φ^{∗}(c, −). But the object mapping ℂ_{0} [figure] 𝔸_{0}: c ↦ F c is functorial: since ℂ ≤ Φ^{∗} ⊗_{𝔸} Φ it follows that

ℂ(c′, c) ≤ Φ^{∗}(c′, −) ⊗_{𝔸} Φ(−, c) = 𝔸(F c′, −) ⊗_{𝔸} 𝔸(−, F c) = 𝔸(F c′, F c).

Clearly Φ ⊣ Φ^{∗} converges to the functor F: ℂ [figure] 𝔸: c ↦ F c.

The next proposition explains Cauchy completeness as “admitting all Cauchy-weighted colimits”.

7.2. Proposition. A Q-category is Cauchy complete if and only if it admits all colimits weighted by a Cauchy distributor, if and only if it admits all colimits weighted by a Cauchy (contravariant) presheaf.

Proof. Let 𝔹 be Cauchy complete; given a Cauchy distributor Φ: ℂ [figure] 𝔸 and a functor F: 𝔸 [figure] 𝔹, the Φ-weighted colimit of F is the functor K: ℂ [figure] 𝔹 to which the Cauchy distributor 𝔹(−, F −) ⊗_{𝔸} Φ converges:

𝔹(K−, −) = Φ^{∗} ⊗_{𝔸} 𝔹(F −, −) = [Φ, 𝔹(F −, −)].

Suppose now that 𝔹 admits all colimits weighted by a Cauchy presheaf; then every such Cauchy presheaf is representable because ϕ: ∗_{C} [figure] 𝔹 (with ϕ ⊣ ϕ^{∗}) converges to colim(ϕ, 1_{𝔹}):

𝔹( colim(ϕ, 1_{𝔹})−, −) = [ϕ, 𝔹(1_{𝔹}−, −)] = ϕ^{∗} ⊗_{𝔹} 𝔹(−, −) = ϕ^{∗}.

By 7.1 the category 𝔹 is Cauchy complete. (The remaining implication is trivial.)

‘Cauchy completeness’ is a self-dual notion: a Q-category 𝔸 is Cauchy complete if and only if the Q^{op}-category 𝔸^{op} is. (This really follows directly from the definition.) Therefore we have – by dualizing the statement of 7.2, using that ‘Cauchy completeness’ is a self-dual notion – at once the following.

7.3. Proposition. A Q-category 𝔸 is Cauchy complete if and only if it admits all limits weighted by a right adjoint distributor, if and only if it admits all limits weighted by a right adjoint (covariant) presheaf.

In the rest of this section we work in terms of colimits.

7.4. Example. Every 2-enriched category 𝔸 is Cauchy complete: a down-set of 𝔸 (i.e. a contravariant presheaf on 𝔸) is Cauchy if and only if it is principal (i.e. a representable presheaf).

7.5. Example. The Rel(Ω)-category associated to a sheaf on a locale Ω (see 3.11) is (skeletal and) Cauchy complete. In fact, it is also symmetric in the sense that it equals its opposite (which makes sense because Rel(Ω)^{op} = Rel(Ω)). [Walters, 1981] proves that sheaves on Ω are precisely the symmetric, skeletal, Cauchy complete Rel(Ω)- categories. Dropping symmetry and skeletality, Cauchy complete Rel(Ω)-categories are ordered sheaves on Ω [Betti et al., 1983; Stubbe, 2004c].

7.6. Example. Consider again Lawvere’s generalized metric spaces (see 3.12); denote by 𝔸 such a space. In [Lawvere, 1973] it is proved that a left adjoint distributor from a one-point space into 𝔸 is precisely (an equivalence class of) a Cauchy sequence in 𝔸—and it is therefore that left adjoint distributors are also called Cauchy distributors. Cauchy completeness of 𝔸 as enriched category coincides with its Cauchy completeness as metric space (“all Cauchy sequences/distributors converge”).

Cauchy completing a category

Not every Q-category 𝔸 is Cauchy complete; still we are interested in “classifying” the Cauchy distributors on 𝔸: we want to construct a Q-category 𝔸_{cc} such that

for every Q-category ℂ, Map(Dist(Q))(ℂ, 𝔸) ≃ Cat(Q)(ℂ, 𝔸_{cc}).  (14)

Much like the construction of P𝔸 we have the following proposition (the proof of which is much like that of 6.1, so it is omitted).

7.7. Proposition. For any Q-category 𝔸, the “Cauchy completion 𝔸_{cc} of 𝔸” is, by definition, the full subcategory of P𝔸 determined by the Cauchy presheaves on 𝔸:

(𝔸_{cc})_{0} = {ϕ ∈ (P𝔸)_{0} | ϕ is Cauchy}.

This Q-category is skeletal and it satisfies (14) in the following way:

- for a Cauchy distributor Φ: ℂ [figure] 𝔸 the corresponding functor F_{Φ}: ℂ [figure] 𝔸_{cc} maps c ∈ ℂ_{0} onto the Cauchy presheaf Φ(−, c): tc [figure] 𝔸;

- and for a functor F: ℂ [figure] 𝔸_{cc} the corresponding Cauchy distributor Φ_{F}: ℂ [figure] 𝔸 has elements, for (c, a) ∈ ℂ_{0}×𝔸_{0}, Φ_{F} (a, c) = F (c)(a) (and its right adjoint has elements Φ_{F}^{∗} (c, a) = F (c)^{∗}(a)).

Note that the hom-arrows in 𝔸_{cc} are identical to those in P𝔸, but since objects ϕ, ψ ∈ 𝔸_{cc} have right adjoints in Dist(Q) we may write (with slight abuse of notation) that 𝔸_{cc}(ψ, ϕ) = ψ^{∗} ⊗_{𝔸} ϕ. This will be a helpful trick in many of the calculations that follow.

Plugging the identity distributor 𝔸: 𝔸 [figure] 𝔸 into the equivalence (14) gives us a functor

i_{𝔸}: 𝔸 [figure] 𝔸_{cc}: a ↦ (𝔸(−, a): ∗_{ta} [figure] 𝔸)

which is precisely the factorization of the Yoneda embedding Y_{𝔸}: 𝔸 [figure] P𝔸 over the full inclusion j_{𝔸}: 𝔸_{cc} [figure] P𝔸 that defined 𝔸_{cc}. An alternative for the classifying property in (14) is then: for any ℂ, a distributor Φ: ℂ [figure] 𝔸 is Cauchy if and only if the corresponding functor F_{Φ}: ℂ [figure] P𝔸 (as in 6.1) factors through j_{𝔸}: 𝔸_{cc} [figure] P𝔸. It follows too that i_{𝔸} enjoys properties similar to those of Y_{𝔸} (especially 6.3 and 6.5).

7.8. Proposition. For any Cauchy presheaf ϕ: ∗_{C} [figure] 𝔸 and any object a ∈ 𝔸_{0}, 𝔸_{cc}(i_{𝔸}(a), ϕ) = ϕ(a) and 𝔸_{cc}(ϕ, i_{𝔸}(a)) = ϕ^{∗}(a). In particular is i_{𝔸}: 𝔸 [figure] 𝔸_{cc} fully faithful.

7.9. Proposition. Every Cauchy presheaf on a Q-category 𝔸 is canonically the colimit of representables in 𝔸_{cc}: for ϕ: ∗_{C} [figure] 𝔸 Cauchy, the ϕ-weighted colimit of i_{𝔸}: 𝔸 [figure] 𝔸_{cc} (exists and) is ϕ itself. In other words, colim(ϕ, i_{𝔸}) = ϕ (where we may write an equality instead of merely an isomorphism, for 𝔸_{cc} is skeletal).

Up to now we have discussed many similarities between Y_{𝔸}: 𝔸 [figure] P𝔸 and i_{𝔸}: 𝔸 [figure] 𝔸_{cc}; but here are two remarkable differences.

7.10. Proposition. A Q-category 𝔸 is Cauchy complete if and only if i_{𝔸}: 𝔸 [figure] 𝔸_{cc} is an essentially surjective functor, if and only if it is an equivalence in Cat(Q)^{12}.

Proof. Since i_{𝔸}: 𝔸 [figure] 𝔸_{cc} is fully faithful, it is clear that it is an equivalence if and only if it is essentially surjective. But the essential surjectivity, saying that any ϕ ∈ 𝔸_{cc} is isomorphic to i_{𝔸}(a) for some a ∈ 𝔸, is precisely the statement in 7.1.

7.11. Proposition. Any Q-category 𝔸 is isomorphic to its Cauchy completion in the quantaloid Dist(Q)^{13}.

Proof. i_{𝔸}: 𝔸 [figure] 𝔸_{cc} is a fully faithful functor, so the unit of the induced adjoint pair of distributors 𝔸_{cc}(−, i_{𝔸}−) ⊣ 𝔸_{cc}(i_{𝔸}−, −) is an equality. It suffices to prove that also its co-unit is an equality. Thereto, consider ϕ, ψ ∈ (𝔸_{cc})_{0} and calculate that

𝔸_{cc}(ψ, i_{𝔸}−) ⊗_{𝔸} 𝔸_{cc}(i_{𝔸}−, ϕ) = ⋁_{a∈𝔸_{0}} 𝔸_{cc}(ψ, 𝔸(−, a)) ∘ 𝔸_{cc}(𝔸(−, a), ϕ)
= ψ^{∗} ⊗_{𝔸} ⋁_{a∈𝔸_{0}} (𝔸(−, a) ∘ 𝔸(a, −))⊗_{𝔸}ϕ
= ψ^{∗} ⊗_{𝔸} 𝔸 ⊗_{𝔸} ϕ
= ψ^{∗} ⊗_{𝔸} ϕ
= 𝔸_{cc}(ψ, ϕ).

That is to say, 𝔸_{cc}(−, i_{𝔸}−) ⊗_{𝔸} 𝔸_{cc}(i_{𝔸}−, −) = 𝔸_{cc}, as wanted.

The following result, although it uses 7.11 in its proof, may be considered analogous to 6.4.

7.12. Proposition. For any 𝔸, the Cauchy completion 𝔸_{cc} is Cauchy complete.

Proof. It suffices to prove that any Cauchy presheaf on 𝔸_{cc} is representable: for any Cauchy Φ: ∗_{C} [figure] 𝔸_{cc} we must find a Cauchy ϕ: ∗_{C} [figure] 𝔸 such that 𝔸_{cc}(−, ϕ) = Φ. Using 𝔸 ≅ 𝔸_{cc} in Dist(Q) (see 7.11) we put ϕ := 𝔸_{cc}(i_{𝔸}−, −) ⊗_{𝔸_{cc}} Φ,

[figure: ∗_{C}; Φ; ϕ; 𝔸_{cc}; 𝔸_{cc}(i_{𝔸}−, −); 𝔸,]

12Cocompleteness of 𝔸 is in general not equivalent to Y_{𝔸}: 𝔸 [figure] P𝔸 being an equivalence in Cat(Q)!

13But 𝔸 need not be isomorphic to P𝔸 in Dist(Q)!

which is Cauchy because it is the composition of Cauchy distributors (𝔸_{cc}(i_{𝔸}−, −) is not only right but also left adjoint because it is invertible). Now for any ψ ∈ 𝔹_{cc},

𝔸_{cc}(ψ, ϕ) = ψ^{∗} ⊗_{𝔸} ϕ
= ψ^{∗} ⊗_{𝔸} 𝔸_{cc}(i_{𝔸}−, −) ⊗_{𝔸_{cc}} Φ
= 𝔸_{cc}(ψ, i_{𝔸}−) ⊗_{𝔸} 𝔸_{cc}(i_{𝔸}−, −) ⊗_{𝔸_{cc}} Φ
= 𝔸_{cc}(ψ, −) ⊗_{𝔸_{cc}} Φ
= Φ(ψ),

so ϕ indeed represents Φ.

A direct consequence of 7.11 and 7.12 is that Dist(Q) is equivalent to its full subquantaloid Dist_{cc}(Q) determined by the Cauchy complete categories.

In view of the examples, in particular 7.5, it is tempting to think of the Cauchy completion 𝔸 ↦ 𝔸_{cc} of a Q-category as a sheafification. The following will only strengthen that intuition.

The universality of the Cauchy completion

The proof of the following is analogous to that of 6.9, and therefore it is omitted.

7.13. Proposition. A functor F: 𝔸 [figure] 𝔹 has a pointwise left Kan extension along i_{𝔸}: 𝔸 [figure] 𝔸_{cc} if and only if, for every ϕ ∈ 𝔸_{cc}, the ϕ-weighted colimit of F exists; in this case ⟨F, i_{𝔸}⟩∘i_{𝔸} ≅ F^{14}. Actually, ⟨F, i_{𝔸}⟩ is the essentially unique factorization of F through i_{𝔸}.

In other words, denoting Cat_{cc}(Q) for the full sub-2-category of Cat(Q) determined by the Cauchy complete categories, we have for Q-categories 𝔸 and 𝔹, with 𝔹 Cauchy complete, an equivalence induced by composition with i_{𝔸}: 𝔸 [figure] 𝔸_{cc}

Cat_{cc}(Q)(𝔸_{cc}, 𝔹) [figure: ∼] Cat(Q)(𝔸, 𝔹): G ↦ G ∘ i_{𝔸}

(with inverse F ↦ ⟨F, i_{𝔸}⟩).

7.14. Proposition. Cauchy completing provides a left (bi)adjoint to the embedding of Cat_{cc}(Q) in Cat(Q): (−)_{cc} ⊣ U: Cat_{cc}(Q) [figure] Cat(Q).

The 2-functor

Cat_{cc}(Q) [figure] Map(Dist_{cc}(Q)): (F: 𝔸 [figure] 𝔹) ↦ (𝔹(−, F −): 𝔸 [figure] 𝔹)

is the identity on objects, and induces equivalences of the hom-orders: it is a biequivalence. Since Dist(Q) itself is equivalent to Dist_{cc}(Q), we may record that

Cat_{cc}(Q) ≃ Map(Dist_{cc}(Q)) ≃ Map(Dist(Q)).

Finally we wish to discuss the “interplay” between 𝔸, P𝔸, P^{†}𝔸 and 𝔸_{cc}.

14But it is not true in general that ⟨F, i_{𝔸}⟩: 𝔸_{cc} [figure] 𝔹 is a left adjoint!

7.15. Proposition. For Q-categories 𝔸 and 𝔹, the following are equivalent:

1. 𝔸 ≅ 𝔹 in Dist(Q);

2. 𝔸_{cc} ≃ 𝔹_{cc} in Cat(Q);

3. P𝔸 ≃ P𝔹 in Cat(Q);

4. P^{†}𝔸 ≃ P^{†}𝔹 in Cat(Q).

Proof. (1 ⇔ 2) 𝔸 ≅ 𝔹 in Dist(Q) implies and is implied by 𝔸_{cc} ≅ 𝔹_{cc} in Dist_{cc}(Q) (using 7.7), which by the biequivalence Cat_{cc}(Q) ≃ Map(Dist_{cc}(Q)) means that 𝔸_{cc} ≃ 𝔹_{cc} in Cat_{cc}(Q), or in Cat(Q).

(1 ⇔ 3) 𝔸 ≅ 𝔹 in Dist(Q) if and only if P𝔸 ≃ P𝔹 in Cocont(Q) by the local equivalence Dist(Q) [figure] Cocont(Q) of 6.12; but to give an equivalence between cocomplete categories in Cocont(Q) is the same as in Cat(Q) (because the functors constituting the equivalence are always left adjoint, thus cocontinuous).

(1 ⇔ 4) is dual to the previous argument.

Whenever for two Q-categories 𝔸 and 𝔹 the equivalent conditions of 7.15 are fulfilled, then 𝔸 is Morita equivalent to 𝔹. The morality is then that “a Q-category is but a presentation of its Cauchy completion”; in particular, if one only cares about Q-category theory “up to Cauchy completion” (i.e. “up to Morita equivalence”), then one can forget about functors and Cauchy completion and Morita equivalence altogether, and treat Map(Dist(Q)) as “the 2-category of Q-categories and functors”.

8. Appendix: Distributor calculus

Lax (co)limits in a quantaloid

Write Q for a quantaloid, and D for a small category. A lax functor F: D [figure] Q is a mapping of objects D_{0} [figure] Q_{0}: D ↦ F D together with actions D(D, D′) [figure] Q(F D, F D′): f ↦ F f on hom-sets such that

- for all D ∈ D_{0}, 1_{F D} ≤ F 1_{D};
- for all composable f, g ∈ D, (F g ∘ F f) ≤ F (g ∘ f).

Given two such lax functors F, G: D [figure] Q, a family (θ_{D}: F D [figure] GD)_{D∈D_{0}} of Q-arrows constitutes a lax-natural transformation θ: F [figure] G if

- for all f: D [figure] D′ in D, (θ_{D′} ∘ F f) ≥ (Gf ∘ θ_{D});

and the family is said to constitute an oplax-natural transformation if

- for all f: D [figure] D′ in D, (θ_{D′} ∘ F f) ≤ (Gf ∘ θ_{D}).

The collection of lax transfos, resp. oplax transfos, between two parallel lax functors F, G: D [figure] Q constitutes a sup-lattice (with “componentwise supremum”); we shall denote it by Lax(F, G), resp. OpLax(F, G). Lax transfos (oplax transfos) may also be composed: this too is done “componentwise”. The notions of ‘lax transfo’ and ‘oplax transfo’ are dual in the following sense: two lax functors F, G: D [figure] Q may be thought of as F^{op}, G^{op}: D^{op} [figure] Q^{op}, and then we have that Lax(F, G) = OpLax(G^{op}, F^{op}).

Any (ordinary) functor F: D [figure] Q is trivially lax, and any (ordinary) natural transformation α: F [figure] G between functors is trivially (op)lax. The set Nat(F, G) of natural transformations between two functors F, G: D [figure] Q is a sub-sup-lattice of both the suplattices Lax(F, G) and OpLax(F, G) (it is their intersection).

In what follows, denote by Δ_{A}: D [figure] Q the constant functor at the object A ∈ Q (on whatever category D). For an arrow k: A [figure] B in Q the constant transformation from Δ_{A} to Δ_{B} will be denoted by Δ_{k}.

8.1. Definition. Let D be a small category, and Q any quantaloid. For a lax functor F: D [figure] Q a lax cone over F consists of an object L ∈ Q and a lax transfo π ∈ Lax(Δ_{L}, F). Such a lax cone (L, π) is the (necessarily essentially unique) lax conical limit of F if it is universal: for all A ∈ Q there is an isomorphism^{15} of sup-lattices induced by composition with π: Δ_{L} [figure] F,

Q(A, L) [figure: ∼] Lax(Δ_{A}, F): k ↦ π ∘ Δ_{k}.  (15)

Lax cocone and lax conical colimit are the dual notions: a lax cocone on F is an oplax transfo σ ∈ OpLax(F, Δ_{C}), and it is the lax colimit if it is universal.

So a lax functor F: D [figure] Q admits a lax limit (lax colimit) in Q if and only if F^{op}: D^{op} [figure] Q^{op} admits a lax colimit (lax limit) in Q^{op}—this is the duality between limits and colimits.

In general, for an ordinary functor F: D [figure] Q, the limit on F is not necessarily its lax limit: whereas any cone α ∈ Nat(Δ_{L}, F) is trivially also a lax cone, it is not true that the universal property for such an α to be a limit implies the universal property for α to be a lax limit. But in some cases lax limits do coincide with ordinary ones. (And by duality we can apply the result to lax colimits, of course.)

8.2. Proposition. Suppose that, for some lax functor F_{1}: D_{1} [figure] Q and some (ordinary) functor F_{2}: D_{2} [figure] Q into a quantaloid Q, for all X ∈ Q there is a sup-isomorphism

Lax(Δ_{X}, F_{1}) [figure: ∼] Nat(Δ_{X}, F_{2})  (16)

which is natural in X. Then some π ∈ Lax(Δ_{L}, F_{1}) is the lax limit of F_{1} if and only if π′ ∈ Nat(Δ_{L}, F_{2}), the image of π under the isomorphism above (taking X = L), is the limit of F_{2}.

15Actually the definition of “lax (co)limit” in a general bicategory asks only for an equivalence, and not for an isomorphism. But an equivalence between antisymmetric orders is always an isomorphism, so we put that right away in this definition.

Proof. To say that the functor F_{2}: D_{2} [figure] Q has limit π′ ∈ Nat(Δ_{L}, F_{2}) is to say that there are bijections

β′_{X}: Q(X, L) [figure: ∼] Nat(Δ_{X}, F_{2}): f ↦ π′ ∘ Δ_{f}

defining an isomorphism of the Set-valued functors Q(−, L) and Nat(Δ_{−}, F_{2}) on (the underlying category of) Q^{op} in CAT; and π′ = β′_{L}(1_{L}). However, not only is each Nat(Δ_{X}, F_{2}) is a sup-lattice, Nat(Δ_{−}, F_{2}) also preserves suprema of Q-arrows; it is thus a Sup-valued homomorphism on Q^{op}. Moreover, the natural bijections above – since they are induced by composition in Q – always preserve suprema. Therefore, F_{2} has limit π′ if and only if β′: Q(−, L) [figure] Nat(Δ_{−}, F_{2}) is an isomorphism of Sup-valued homomorphisms on Q^{op} in QUANT.

On the other hand, to say that π ∈ Lax(Δ_{X}, F_{1}) is the lax limit of the lax functor F_{1}: D_{1} [figure] Q, is to say that there are sup-isomorphisms

β_{X}: Q(X, L) [figure: ∼] Lax(Δ_{X}, F_{1}): f ↦ π ∘ Δ_{f}

defining an isomorphism β: Q(−, L) [figure] Lax(Δ_{−}, F_{1}) of Sup-valued homomorphisms on Q^{op} in QUANT; and π = β_{L}(1_{L}).

The result above then follows by composition of isomorphisms of Sup-valued homomorphisms on Q^{op} in QUANT.

Two important examples

A set-indexed family (A_{i})_{i∈I} of objects in a quantaloid Q can be identified with (the image of) the functor F: I [figure] Q : i ↦ A_{i} whose domain is the index-set regarded as discrete category.

8.3. Proposition. For a set-indexed family (A_{i})_{i∈I} of objects in a quantaloid Q, the following are equivalent:

1. (A_{i})_{i∈I} has a limit (i.e. product) in Q;

2. (A_{i})_{i∈I} has a colimit (i.e. coproduct) in Q;

3. (A_{i})_{i∈I} has a lax limit in Q;

4. (A_{i})_{i∈I} has a lax colimit in Q;

5. there exist an object A and arrows

( [figure: A; p_{i}; A_{i}; s_{i}; A] )_{i∈I}

in Q such that, for all i, j ∈ I, p_{j} ∘ s_{i} = δ_{A_{i},A_{j}} and ⋁_{i∈I}(s_{i} ∘ p_{i}) = 1_{A}.

The cone (p_{i}: A [figure] A_{i})_{i∈I} is then the (lax) limit of the given family, and the cocone (s_{i}: A_{i} [figure] A)_{i∈I} is its (lax) colimit. For each i ∈ I the coprojection s_{i}: A_{i} [figure] A is left adjoint to the projection p_{i}: A [figure] A_{i}.

Proof. The equivalence (1 ⇔ 2 ⇔ 5) is well-known. (1 ⇔ 3) and (2 ⇔ 4) follow from 8.2: because the diagram F: I [figure] Q: i ↦ A_{i} is discrete, any lax cone on F is in fact an ordinary cone; and any lax cocone is an ordinary cocone. That is to say, in this particular case we have for any object A ∈ Q that Nat(Δ_{A}, F) = Lax(Δ_{A}, F) and Nat(F, Δ_{A}) = OpLax(F, Δ_{A}). The adjunctions s_{i} ⊣ p_{i} hold since 1_{A_{i}} = p_{i} ∘ s_{i} and s_{i} ∘ p_{i} ≤ 1_{A}.

As usual, we will refer to limits-colimits (in a quantaloid Q) as direct sums and we write them as ⊕_{i}A_{i}.

A monad in a quantaloid Q is an endo-arrow A[figure: t] such that t ∘ t ≤ t and 1_{A} ≤ t. It can be identified with (the image of) a lax functor F: Ob [figure] Q: ∗ ↦ A, 1_{∗} ↦ t. Here Ob stands for the “generic object”: the category with one object ∗ and one arrow 1_{∗}. On the other hand, A[figure: t] is a diagram in Q, thus (the image of) an ordinary functor F′: Ar [figure] Q : ∗_{1} ↦ A, ∗_{2} ↦ A, α ↦ t. Now Ar stands for the “generic arrow”: the category with two objects, ∗_{1} and ∗_{2}, and exactly one arrow that is not an identity, α: ∗_{1} [figure] ∗_{2}.

8.4. Proposition. For a monad A[figure: t]in a quantaloid Q, the following are equivalent:

1. A[figure: t]has a limit (i.e. t, 1_{A}: A [figure] A have an equalizer) in Q;

2. A[figure: t]has a colimit (i.e. t, 1_{A}: A [figure] A have a coequalizer) in Q;

3. A[figure: t]has a lax limit in Q;

4. A[figure: t]has a lax colimit in Q;

5. there exist an object B and arrows

[figure: B; p; A; s; B]

in Q such that p ∘ s = t and s ∘ p = 1_{B}.

If the above is the case for the given monad t, then p: B [figure] A is actually its (lax) limit, and s: A [figure] B its (lax) colimit. And the coprojection s: A [figure] B is left adjoint to the projection p: B [figure] A.

Proof. (1 ⇔ 2 ⇔ 5) is well-known for idempotents in a category, so in particular applies to monads in a quantaloid (it is immediate that a monad in Q is an idempotent). (1 ⇔ 3) and (2 ⇔ 4) follow from 8.2: because 1_{A} ≤ t, it is the same to give a lax cone on F: Ob [figure] Q as to give an ordinary cone on F′: Ar [figure] Q; and lax cocones on F coincide with cocones on F′. More precisely, we have for any object X ∈ Q identical sup-lattices Nat(Δ_{X}, F′) = Lax(Δ_{X}, F) and Nat(F′, Δ_{X}) = OpLax(F, Δ_{X}). The adjunction s ⊣ p follows because 1_{A} ≤ t = p ∘ s and s ∘ p = 1_{B}.

The lax limit of a monad A[figure: t]in a 2-category is called the object of t-algebras, or the Eilenberg-Moore object for t, and often written as A^{t}. And the lax colimit is the object of free t-algebras, or the Kleisli object, written A_{t}. In a quantaloid Q, A^{t} is isomorphic to A_{t} as soon as one (and thus also the other) exists: “every t-algebra is free”! When Q admits (free) algebra objects for all of its monads, we will say that all monads in Q split.

Universal property of the distributor calculus

For the following we refer also to [Betti et al., 1983; Carboni et al., 1987].

Given a quantaloid Q, there is a quantaloid Matr(Q) of “matrices with elements in Q”, with:

- objects: Q_{0}-typed sets^{16};

- arrows: for two objects (X, t), (Y, t), an arrow 𝕄: (X, t) [figure] (Y, t) is a matrix of Q-arrows 𝕄(y, x): tx [figure] ty, one for each (x, y) ∈ X × Y;

- two-cells: for 𝕄, ℕ: (X, t) [figure] (Y, t), 𝕄 ≤ ℕ when 𝕄(y, x) ≤ ℕ(y, x) for all x, y;

- composition: for 𝕄: (X, t) [figure] (Y, t) and ℕ: (Y, t) [figure] (Z, t), the elements of the composite matrix ℕ ∘ 𝕄: (X, t) [figure] (Y, t) are (ℕ ∘ 𝕄)(z, x) = ⋁_{y∈Y} ℕ(z, y) ∘ 𝕄(y, x);

- identities: the identity matrix 𝕀_{(X,t)}: (X, t) [figure] (Y, t) has elements 𝕀_{(X,t)}(x′, x) = δ_{tx,tx′} (“Kronecker deltas”).

Clearly, Q can be embedded in Matr(Q): every arrow f: A [figure] B in Q may be viewed as a one-element matrix. We write this embedding as i_{Q}: Q [figure] Matr(Q). Obviously, i_{Q}: Q [figure] Matr(Q) is a fully faithful homomorphism of quantaloids; it is an equivalence of quantaloids if and only if Q has all direct sums. Actually, Matr(Q) itself has all direct sums (so “taking matrices” is an essentially idempotent process), and i_{Q}: Q [figure] Matr(Q) describes the universal direct-sum completion of Q in QUANT.

Given a quantaloid Q, the following defines a quantaloid Mnd(Q) of “monads and bimodules in Q”:

- objects: monads in Q;

- arrows: an arrow b: t [figure] s between monads A[figure: t]and B[figure: s]is a bimodule from t to s, i.e. a Q-arrow b: A [figure] B such that s ∘ b ≤ b and b ∘ t ≤ b;

- two-cells: put b ≤ c in Mnd(Q)(t, s) whenever b ≤ c in Q(A, B);

- composition: the composition c ⊗_{s} b: t [figure] r of two bimodules b: t [figure] s and c: s [figure] r is c ⊗_{s} b = c ∘ b in Q;

16Recall that such a Q_{0}-typed set (X, t) is really just a way of writing a small discrete diagram (tx)_{x∈X} in Q.

- identity: on a monad A[figure: t]the identity bimodule is t: t [figure] t itself.

The “tensor” notation for composition of bimodules reminds us that the identity on a monad A[figure: t]is not 1_{A} but rather t itself.

An arrow f: A [figure] B may be regarded as bimodule from 1_{A} to 1_{B}; denote the corresponding embedding of quantaloids by j_{Q}: Q [figure] Mnd(Q). Then j_{Q}: Q [figure] Mnd(Q) is a fully faithful homomorphism of quantaloids which is an equivalence of quantaloids if and only if in Q all monads split. In the quantaloid Mnd(Q) all monads split (so “taking monads and bimodules” is an essentially idempotent process), actually, j_{Q}: Q [figure] Mnd(Q) describes the universal split-monad completion of Q in QUANT.

It is now a matter of fact that, for any quantaloid Q, Dist(Q) = Mnd(Matr(Q)). Indeed, take for example the definition of “Q-category”: such an 𝔸 consists of an endomatrix of hom-arrows 𝔸(a′, a): ta [figure] ta′ on the Q_{0}-typed set (𝔸_{0}, t) of objects; these data ought to satisfy, for all a, a′, a′′ ∈ 𝔸_{0},

𝔸(a′′, a′) ∘ 𝔸(a′, a) ≤ 𝔸(a′′, a) and 1_{ta} ≤ 𝔸(a, a),

or equivalently,

⋁_{a′∈𝔸_{0}} 𝔸(a′′, a′) ∘ 𝔸(a′, a) ≤ 𝔸(a′′, a) and δ_{ta,ta′} ≤ 𝔸(a′, a),

or still equivalently, but now in terms of matrices, 𝔸 ∘ 𝔸 ≤ 𝔸 and 𝕀_{𝔸_{0}} ≤ 𝔸. This says precisely that 𝔸 is a monad in Matr(Q), thus an object of Mnd(Matr(Q)). (Similar for distributors.)

Consequently we now know the universal property of Dist(Q): the composite embedding k_{Q}: Q [figure] Dist(Q) given by j_{Matr(Q)}∘i_{Q} is a fully faithful homomorphism of quantaloids which is an equivalence if and only if Q has all direct sums and all monads split. Dist(Q) itself has all direct sums and all monads split (“taking categories and distributors” is an essentially idempotent process); actually, k_{Q}: Q [figure] Dist(Q) describes the universal directsum-and-split-monad completion of Q in QUANT.

Having all direct sums and a splitting for all monads, is enough for a quantaloid Q to admit all lax (co)limits.

8.5. Proposition. For any quantaloid Q, the following are equivalent:

1. Q has direct sums and all monads split;

2. Q has all lax limits;

3. Q has all lax colimits;

4. for any lax functor F: D [figure] Q there exist an object L and arrows

( [figure: L; p_{D}; F D; s_{D}; L] )_{D∈D_{0}}

in Q such that

⋁{s_{D} ∘ p_{D} | D ∈ D_{0}} = 1_{L} and p_{D′} ∘ s_{D} = ⋁{F d | d ∈ D(D, D′)}.

With notations as in the last sentence, the (p_{D}: L [figure] F D)_{D∈D_{0}} form the lax limit of F, and the (s_{D}: F D [figure] L)_{D∈D_{0}} its lax colimit; moreover, for every D ∈ D_{0}, the coprojection s_{D} is left adjoint to the projection p_{D}.

Proof. (2 ⇒ 1) and (3 ⇒ 1) follow from 8.3 and 8.4.

For (4 ⇒ 2) and (4 ⇒ 3), first observe that s_{D} ⊣ p_{D}: on the one hand 1_{L} = ⋁{s_{D} ∘p_{D} | D ∈ D_{0}} ≥ s_{D} ∘ p_{D}, and on the other p_{D} ∘ s_{D} = ⋁{F d | d ∈ D(D, D)} ≥ F (1_{D}) ≥ 1_{F D}. Knowing this, it follows from p_{D′} ∘ s_{D} = ⋁{F d | d ∈ D(D, D′)} that for every d: D [figure] D′, p_{D′} ∘ s_{D} ≥ F d, hence also s_{D} ≥ s_{D′} ∘ F d and p_{D′} ≥ F d ∘ p_{D}. Hence π = (p_{D})_{D} is a lax cone, and σ = (s_{D})_{D} a lax cocone, on F. As for their universality, we must verify that for any X ∈ Q the sup-morphisms

Q(X, L) [figure] Lax(Δ_{X}, F): x ↦ π ∘ Δ_{x},
Q(L, X) [figure] Lax(F, Δ_{X}): x ↦ Δ_{x} ∘ σ

are bijective (in which case they are sup-isomorphisms). But a calculation shows that their respective inverses are

π′ = (p′_{D})_{D} ↦ ⋁{s_{D} ∘ p′_{D} | D ∈ D_{0}},
σ′ = (s′_{D})_{D} ↦ ⋁{s′_{D} ∘ p_{D} | D ∈ D_{0}}.

For (1 ⇒ 4), note that a lax functor F: D [figure] Q determines a Q-category 𝔻 with objects 𝔻_{0} = D_{0}, types tD = F D, and hom-arrows 𝔻(D′, D) = ⋁{F d | d ∈ D(D, D′)}: the inequalities 𝔻(D′′, D′) ∘ 𝔻(D′, D) ≤ 𝔻(D′′, D) and 1_{F D} ≤ 𝔻(D, D) hold precisely because F is lax! Since Q has all direct sums and all monads split, this Q-category 𝔻 must be isomorphic in Dist(Q) to a one-object Q-category, call it ∗_{L}. This means that there exist distributors σ: 𝔻 [figure] ∗_{L} and π: ∗_{L} [figure] 𝔻 such that σ ⊗_{∗_{L}} π = 𝔻 and π ⊗_{𝔻}σ = 1_{L}. Such σ and π are really collections of Q-arrows

( [figure: L; p_{D} := π(D); F D; s_{D} := σ(D); L] )_{D∈D_{0}}

satisfying

p_{D′} ∘ s_{D} = 𝔻(D′, D) = ⋁{F d | d ∈ D(D, D′)},
⋁{s_{D} ∘ p_{D} | D ∈ 𝔻_{0}} = ⋁{s_{D} ∘ p_{D} | D ∈ D_{0}} = 1_{L}.

44  ISAR STUBBE

References

[1] Jean Bénabou (1963) Catégories avec multiplication, C. R. Acad. Sci. Paris 256, pp. 1887–1890.

[2] Jean Bénabou (1965) Catégories relatives, C. R. Acad. Sci. Paris 260, pp. 3824–3827.

[3] Jean Bénabou (1967) Introduction to bicategories, Lecture Notes in Math. 47, pp. 1–77.

[4] Renato Betti, Aurelio Carboni, Ross H. Street and Robert F. C. Walters (1983) Variation through enrichment, J. Pure Appl. Algebra 29, pp. 109–127.

[5] Francis Borceux and Rosanna Cruciani (1998) Skew Ω-sets coincide with Ω-posets, Cahiers Topol. Géom. Différ. Catég. 39, pp. 205–220.

[6] Aurelio Carboni, Stefano Kasangian and Robert F. C. Walters (1987) An axiomatics for bicategories of modules, J. Pure Appl. Algebra 45, pp. 127–141.

[7] Samuel Eilenberg and G. Max Kelly (1966) Closed categories, pp. 421–562 in Proc. Conf. Categorical Algebra (La Jolla 1965), Springer, New York.

[8] Robert Gordon and A. John Power (1997) Enrichment through variation, J. Pure Appl. Algebra 120, pp. 167–185.

[9] G. Max Kelly (1982) Basic concepts of enriched category theory. London Mathematical Society Lecture Note Series. Cambridge University Press, Cambridge.

[10] Anna Labella and Vincent Schmitt (2002) Change of base, Cauchy completeness and reversibility, Theory Appl. Categ. 10, pp. 187–219.

[11] F. William Lawvere (1973) Metric spaces, generalized logic and closed categories, Rend. Sem. Mat. Fis. Milano 43, pp. 135–166. Also in: Reprints in Theory and Applications of Categories, No. 1 (2002) pp. 1–37

[12] Marie-Anne Moens, Ugo Berni-Canani and Francis Borceux (2002) On regular presheaves and regular semicategories, Cahiers Topol. Géom. Différ. Catég. 43, pp. 163–190.

[13] Christopher J. Mulvey (1986) &, Suppl. Rend. Circ. Mat. Palermo II 12, pp. 99–104.

[14] Kimmo I. Rosenthal (1996) The theory of quantaloids. Pitman Research Notes in Mathematics Series. Longman, Harlow.

[15] Ross H. Street (1972) The formal theory of monads, J. Pure Appl. Algebra 2, pp. 149–168.

[16] Ross H. Street (1981) Cauchy characterization of enriched categories, Rend. Sem. Mat. Fis. Milano 51, pp. 217–233. Also in: Reprints in Theory and Applications of Categories, No. 4 (2004) pp. 1–16

[17] Ross H. Street (1983a) Enriched categories and cohomology, Questiones Math. 6, pp. 265–283.

[18] Ross H. Street (1983b) Absolute colimits in enriched categories, Cahiers Topol. Géom. Différ. Catég. 24, pp. 377–379.

[19] Isar Stubbe (2004a) Categorical structures enriched in a quantaloid: tensored and cotensored

categories, preprint, arXiv:math.CT/0411366.

CATEGORICAL STRUCTURES ENRICHED IN A QUANTALOID  45

[20] Isar Stubbe (2004b) Categorical structures enriched in a quantaloid: regular presheaves, regular semicategories, Cahiers Topol. Géom. Différ. Catég. (to appear), arXiv:math.CT/0409475.

[21] Isar Stubbe (2004c) Categorical structures enriched in a quantaloid: orders and ideals over a base quantaloid, Appl. Categ. Structures (to appear), arXiv:math.CT/0409477.

[22] Gilberte Van den Bossche (1995) Quantaloids and non-commutative ring representations, Appl. Categ. Structures 3, pp. 305–320.

[23] Robert F. C. Walters (1981) Sheaves and Cauchy-complete categories, Cahiers Topol. Géom. Différ. Catég. 22, pp. 283–286.

Isar Stubbe Département de Mathématique Université de Louvain Chemin du Cyclotron 2 1348 Louvain-la-Neuve Belgique

Email: i.stubbe@math.ucl.ac.be

This article may be accessed via WWW at http://www.tac.mta.ca/tac/ or by anony-

mous ftp at ftp://ftp.tac.mta.ca/pub/tac/html/volumes/14/1/14-01.{dvi,ps}

THEORY AND APPLICATIONS OF CATEGORIES (ISSN 1201-561X) will disseminate articles that significantly advance the study of categorical algebra or methods, or that make significant new contributions to mathematical science using categorical methods. The scope of the journal includes: all areas of pure category theory, including higher dimensional categories; applications of category theory to algebra, geometry and topology and other areas of mathematics; applications of category theory to computer science, physics and other mathematical sciences; contributions to scientific knowledge that make use of categorical methods.

Articles appearing in the journal have been carefully and critically refereed under the responsibility of members of the Editorial Board. Only papers judged to be both significant and excellent are accepted for publication.

The method of distribution of the journal is via the Internet tools WWW/ftp. The journal is archived electronically and in printed paper format.

Subscription information. Individual subscribers receive (by e-mail) abstracts of articles as they are published. Full text of published articles is available in .dvi, Postscript and PDF. Details will be e-mailed to new subscribers. To subscribe, send e-mail to tac@mta.ca including a full name and postal address. For institutional subscription, send enquiries to the Managing Editor, Robert Rosebrugh,

rrosebrugh@mta.ca.

Information for authors. The typesetting language of the journal is TEX, and L^{A}TEX2e is the preferred flavour. TEX source of articles for publication should be submitted by e-mail directly to an appropriate Editor. They are listed below. Please obtain detailed information on submission format and style files from the journal’s WWW server at http://www.tac.mta.ca/tac/. You may also write to tac@mta.ca to receive details by e-mail.

Managing editor. Robert Rosebrugh, Mount Allison University: rrosebrugh@mta.ca

TEXnical editor. Michael Barr, McGill University: barr@barrs.org

Transmitting editors.

Lawrence Breen, Université Paris 13: breen@math.univ-paris13.fr Ronald Brown, University of North Wales: r.brown@bangor.ac.uk Jean-Luc Brylinski, Pennsylvania State University: jlb@math.psu.edu Aurelio Carboni, Università dell Insubria: aurelio.carboni@uninsubria.it Valeria de Paiva, Xerox Palo Alto Research Center: paiva@parc.xerox.com Martin Hyland, University of Cambridge: M.Hyland@dpmms.cam.ac.uk P. T. Johnstone, University of Cambridge: ptj@dpmms.cam.ac.uk G. Max Kelly, University of Sydney: maxk@maths.usyd.edu.au Anders Kock, University of Aarhus: kock@imf.au.dk Stephen Lack, University of Western Sydney: s.lack@uws.edu.au F. William Lawvere, State University of New York at Buffalo: wlawvere@acsu.buffalo.edu Jean-Louis Loday, Université de Strasbourg: loday@math.u-strasbg.fr Ieke Moerdijk, University of Utrecht: moerdijk@math.uu.nl Susan Niefield, Union College: niefiels@union.edu Robert Paré, Dalhousie University: pare@mathstat.dal.ca Jiri Rosicky, Masaryk University: rosicky@math.muni.cz James Stasheff, University of North Carolina: jds@math.unc.edu Ross Street, Macquarie University: street@math.mq.edu.au Walter Tholen, York University: tholen@mathstat.yorku.ca Myles Tierney, Rutgers University: tierney@math.rutgers.edu Robert F. C. Walters, University of Insubria: robert.walters@uninsubria.it R. J. Wood, Dalhousie University: rjwood@mathstat.dal.ca